### v3

# A Survey on the Safety and Security Threats of Computer-Using Agents

Abstract
Recently, AI-driven interactions with computing devices have advanced from basic prototype tools to sophisticated, LLM-based systems that emulate human-like operations in graphical user interfaces. We are now witnessing the emergence of Computer-Using Agents (CUAs), capable of autonomously performing tasks such as navigating desktop applications, web pages, and mobile apps. However, as these agents grow in capability, they also introduce novel safety and security risks. Vulnerabilities in LLM-driven reasoning, with the added complexity of integrating multiple software components and multimodal inputs, further complicate the security landscape. In this paper, we present a systematization of knowledge on the safety and security threats of CUAs. We conduct a comprehensive literature review and distill our findings along four research objectives: (i) define the CUA that suits safety analysis; (ii) categorize current safety threats among CUAs; (iii) propose a comprehensive taxonomy of existing defensive strategies; (iv) summarize prevailing benchmarks, datasets, and evaluation metrics used to assess the safety and performance of CUAs. Building on these insights, our work provides future researchers with a structured foundation for exploring unexplored vulnerabilities and offers practitioners actionable guidance in designing and deploying secure Computer-Using Agents.
[…]
[Uncaptioned image]JARVIS or [Uncaptioned image]Ultron?
A Survey on the Safety and Security Threats of Computer-Using Agents
Ada Chen1thanks:   Ada Chen, Yongjiang Wu and Junyuan Zhang contribute equally to this paper.  Yongjiang Wu2∗  Junyuan Zhang2∗  Jingyu Xiao2 Shu Yang3  Jen-tse Huang4  Kun Wang5  Wenxuan Wang6thanks:   Wenxuan Wang (jwxwang@gmail.com) is the corresponding authors.  Shuai Wang6 1Carnegie Mellon University  2The Chinese University of Hong Kong 3KAUST  4Johns Hopkins University  5Nanyang Technological University 6The Hong Kong University of Science and Technology
1 Introduction
Large Language Models (LLMs) have evolved rapidly from basic conversational agents to executing complex tasks in diverse computing environments. In particular, Computer-Using Agents (CUAs) have garnered increasing attention and widespread adoption, thanks to their ability to interact with graphical user interfaces (GUIs) in a manner akin to human users (OpenAI, 2025a). Recent systems such as AppAgent, SeeAct, PC-Agent, as well as OpenAI’s o3, and o4-mini introduced in April 2025, highlight the remarkable progress of CUAs (Zhang et al., 2023; Zheng et al., 2024; Liu et al., 2025b; OpenAI, 2025a, b). By integrating multimodal perception, advanced reasoning, and automated control of devices, these agents promise to streamline vast tasks from filling out online forms to executing complex application flows.
Despite the impressive capabilities of CUAs, their operation in real-world settings raises critical safety concerns. Emerging reports reveal that vulnerabilities like visual grounding errors, response delays, and UI interpretation pitfalls can be exploited by malicious attackers, causing unintended or harmful consequences such as data leakage, goal misdirection, and so on (Zheng et al., 2024; Nong et al., 2024; Zhang and Zhang, 2023; Wen et al., 2023; Liu et al., 2025b). Additionally, many of the threats known to standalone LLMs, such as adversarial attacks and jailbreak strategies, now manifest in CUAs with heightened severity, sometimes in new forms adapted to GUI-based environments (Wu et al., 2024a; Kumar et al., 2024; Tian et al., 2023). Novel attack vectors also surface in CUAs, including environment-level manipulations and reasoning-gap attacks that stealthily guide the agent toward risky or undesired behaviors (Wu et al., 2024b; Yuan et al., 2024; Lee et al., 2024a; Zhan et al., 2024). As such, a systematic study on the safety and security threats of CUAs is both timely and necessary.
In this work, we present a comprehensive survey focused on the safety and security threats of Computer-Using Agents (CUAs). First, we propose a unifying definition for CUAs, drawing on a detailed study of state-of-the-art agent systems and workflows. Then, we develop a structured taxonomy of both intrinsic and extrinsic threats by synthesizing literature from the safety of LLM-based agents. After that, we systematically review and categorize existing defense approaches, linking each to the corresponding threat taxonomy. Finally, we summarize various evaluation metrics and datasets for measuring both the severity of threats and the impact of mitigation techniques. Our survey aims to illuminate the landscape of the safety and security study in CUA research to inspire future studies and innovations.
The rest of the paper is organized as follows: Section 2 serves as a background, which defines the concept of a CUA and contextualizes it within existing frameworks. Section 3 details our taxonomy of threats to CUAs, covering both internal vulnerabilities and extrinsic risk factors. Section 4 systematically reviews defense mechanisms and links them to the threat categories they mitigate. Section 5 discusses strategies for systematic evaluation of CUA safety and the effectiveness of defenses. Key insights and highlights are discussed in Section 6. Finally, Section 7 offers concluding remarks and outlines promising directions for future research into safe and robust CUAs.
2 Background
2.1 Computer-Using Agent
In this paper, a Computer-Using Agent (CUA) is an LLM-based system that combines multimodal perception, advanced reasoning, and tool-use capabilities to perceive and interact with graphical user interfaces (GUIs) and external applications just like human users (OpenAI, 2025a). By processing visual information from screenshots, invoking APIs or command-line tools, and executing actions like typing, clicking, and scrolling, a CUA can autonomously perform end-to-end tasks on a computer, such as ordering products, making reservations, and filling out forms (OpenAI, 2025a).
In the realm of agents, several categories fall under the umbrella of Computer-Using Agents:
  • OS Agents: These agents operate within general computing devices, such as desktops and laptops, to perform tasks by interacting with the operating system’s environment and interfaces (Chen et al., 2025d).
  • GUI Agents: Agents that interact specifically with graphical user interfaces to control applications and perform tasks that would typically require human interaction with visual elements (Zhang et al., 2024a).
  • Web Agents: These agents are designed to navigate and interact with web environments, automating tasks such as data retrieval, form submission, and web browsing (Yang et al., 2024a; Liao et al., 2024).
  • Device-control Agents: Agents that manage and control various hardware devices, enabling automation of device-specific operations across different platforms (Zhang and Zhang, 2023; Lee et al., 2024b).
Agent Framework
As an LLM-based agent, the architecture of a CUA comprises the following three core components:
  • Perception: This component enables the agent to gather information from its environment through various input modalities, such as screen reading, system logs, and user inputs.
  • Brain: Serving as the decision-making unit, it processes the information collected by the perception component, interprets it, and formulates appropriate actions with memory mechanisms and planning strategies based on predefined goals and contextual understanding.
  • Action: This component executes the decisions made by the brain, interacting with the operating system, applications, or web interfaces to perform tasks, manipulate data, or control devices as required. Tool use could also be included in this process.
2.2 Literature Review
To organize the studies on the safety and security threats of CUAs, we conducted a comprehensive review of recent literature from 2022 onward. Our search targeted publications that included combinations of the following keywords: GUI Agent, OS Agent, Web Agent, LLM-Based Agent, alongside terms Safety, Security, Attack, Defense, and Threat. Our literature review encompassed several stages:
  1. 1.
    Database Selection: We utilized academic databases and preprint servers, including arXiv, Semantic Scholar, Google Scholar, and OpenReview, to source relevant publications.
  2. 2.
    Keyword Search: After keyword selection, we identified 700+ papers potentially addressing security concerns related to CUAs.
  3. 3.
    Screening and Filtering: Each identified paper underwent a thorough review to assess its relevance. We excluded studies that duplicate or did not directly pertain to security threats or defenses associated with CUAs, resulting in 124 pertinent papers for in-depth analysis.
3 Taxonomy of Safety Threats
3.1 Threat Overview
In this section, we introduce our taxonomy of threats for Computer-Using Agents (CUAs). These threats are categorized into two main types: intrinsic threats and extrinsic threats, which are presented in Table 1 and Table 2, respectively. Intrinsic threats arise from intrinsic aspects of the agent itself, including its training process, configuration, or inherent limitations (Yu et al., 2025; Ferrag et al., 2025). They can induce failures, inefficiencies, or biases in the agent’s functioning, consequently introducing security risks. Extrinsic threats, on the other hand, are initiated by external entities, such as malicious attackers or users, who attempt to exploit vulnerabilities in the agent’s interaction with its surroundings or take advantage of the agent’s intrinsic issues to trigger unsafe behaviors, potentially leading to risky consequences (Yu et al., 2025; Ferrag et al., 2025).
We organize these threats in a tabular format that highlights the following key aspects:
  • Source of the Threats identifies where the threat originates — Environment (Env), Prompt, Model, or User — and indicates whether it serves as a primary contributor ($\blacklozenge$) or a secondary contributor ($\lozenge$) to the threat.
  • Affected Components indicates specific aspects of the agent’s framework (Perception, Brain, and Action) that are vulnerable to potential attacks. A checkmark (✓) shows that a particular component is affected by the threat.
  • Threat Model states the originating entity of each threat.
Threat Source of the Threats Affected Components Threat Model
Env Prompt Model User Perception Brian Action
1 UI Understand&Ground Difficulties $\blacklozenge$ $\checkmark$ Agent Deveploment
2 Scheduling Error $\blacklozenge$ $\checkmark$ Agent Development
3 Misalignment $\blacklozenge$ $\checkmark$ Agent Deployment
4 Hallucination $\lozenge$ $\blacklozenge$ $\checkmark$ Agent Deployment
5 Excessive Context Length $\blacklozenge$ $\checkmark$ Agent Architecture
6 Social and Cultural Concern $\blacklozenge$ $\checkmark$ Agent Training
7 Response Latency $\blacklozenge$ $\checkmark$ $\checkmark$ Deployment / Architecture
8 API Call Error $\blacklozenge$ $\checkmark$ Agent Deployment
Table 1: A taxonomy of intrinsic threats. The symbol $\blacklozenge$ indicates that a threat is fully available to the given item, while $\lozenge$ represents limited availability.
Threat Source of the Threats Affected Components Threat Model
Env Prompt Model User Perception Brian Action
1 Adversarial Attack $\blacklozenge$ $\lozenge$ $\lozenge$ $\checkmark$ Malicious attacker
2 Prompt Injection Attack $\blacklozenge$ $\blacklozenge$ $\lozenge$ $\checkmark$ $\checkmark$ Malicious attacker
3 Jailbreak $\lozenge$ $\blacklozenge$ $\lozenge$ $\checkmark$ $\checkmark$ Malicious attacker
4 Memory Injection Attack $\lozenge$ $\blacklozenge$ $\checkmark$ $\checkmark$ Malicious attacker
5 Backdoor Attack $\lozenge$ $\lozenge$ $\blacklozenge$ $\checkmark$ $\checkmark$ Malicious attacker
6 Reasoning Gap Attack $\lozenge$ $\lozenge$ $\blacklozenge$ $\checkmark$ Malicious attacker
7 System Sabotage $\lozenge$ $\lozenge$ $\blacklozenge$ $\checkmark$ Malicious attacker
8 Web Hacking $\lozenge$ $\lozenge$ $\blacklozenge$ $\checkmark$ Malicious user
Table 2: A taxonomy of extrinsic threats. The symbol $\blacklozenge$ indicates that a threat is fully available to the given item, while $\lozenge$ represents limited availability.
3.2 Intrinsic Threats
Intrinsic threats, which are the issues arising from the agent itself, can lead to a series of negative impacts. In this section, we organize these intrinsic threats, focusing on their mechanisms of action and their corresponding repercussions. Following the overview in Table 1, we discuss each intrinsic threat according to the affected agent framework.
3.2.1 Perception
In the Computer-Using Agents (CUAs), the perception component takes charge of receiving the model input information (e.g. prompt and user instruction), and recognizing the task-specific elements, such as UI screen shots, HTML elements, and other environmental observations. The most common issue in the perception module is the difficulty in UI understanding and grounding.
1
UI Understand and Ground Difficulties  It refers to the challenges faced by models in accurately perceiving, interpreting, and associating UI elements (such as buttons, forms, icons) with semantic meaning, user intent, or external knowledge, due to limitations in layout understanding, semantic ambiguity, or missing contextual grounding.
This challenge stems largely from inherent problems in the available UI datasets. For example, many UI datasets are predominantly static, lacking the dynamic variability seen in real-world applications (Chen et al., 2025c). Additionally, these datasets often suffer from data scarcity, with insufficient samples and task diversity to effectively train models on the wide range of interactions and scenarios encountered in practice (Pahuja et al., 2025). Moreover, the agent sometimes needs to take screenshots controlled at a certain resolution to recognize the current interface, which may lose image details, leading to deficiencies in UI comprehension (Nong et al., 2024).
3.2.2 Brain
The brain component involves reasoning, memory, and planning functions, from which the following six primary threats stem:
2
Scheduling Errors Scheduling errors refer to the internal failures of a CUA agent in managing the execution order, concurrency, or timing of actions, ultimately leading to unintended behaviors. The CUAs needs to handle complex user instructions and interdependent subtasks, and the implementation of the planning function mostly relies on external tools and application-specific APIs to parse the environment into textual elements and interpret predicted actions (Zhang and Zhang, 2023).
Previous studies show that planning before action are essential. In complex tasks, losing the planning has serious negative consequences (Deng et al., 2024). Inaccuracies in task scheduling can disrupt the planned action sequence, leading to inefficiencies and even errors in task execution, which can trigger data leakage and operational privilege issues.
3
Misalignment Misalignment occurs when the agent’s intrinsic reasoning does not properly align with the real-world context or user intent. The problem arises from the pitfalls inherent in LLM. It results in decisions that are out of sync with the environmental demands or user instructions, and potentially unexpected and harmful actions.
Building on this understanding, several studies have explored the underlying causes of misalignment in CUAs. In particular, Ma et al. (2024) highlights that even in benign settings, where both the user and the agent act in good faith and the environment is non-malicious, the presence of unrelated content can distract both generalist and specialist GUI agents, leading to unfaithful behaviors. This observation further underscores the inherent vulnerability of agents to misalignment.
4
Hallucination Hallucination refers to the phenomenon where a CUA agent generates outputs, such as facts, actions, or API calls, that are not grounded in the actual environment, task context, or user input, which primarily stems from insufficient training of agents and their limited grasp of the task-specific knowledge and context.
Among related studies, Mobile-Bench (Deng et al., 2024) highlights that general large models, despite strong reasoning and planning abilities, are prone to generating inaccurate or misleading API calls, revealing a notable form of hallucination within CUAs.
5
Excessive Context Length Excessive context length represents the condition where the accumulated input (e.g., OCR output, HTML, UI trees) to a model, and historical interaction data, exceed or approach the model’s input capacity, leading to degraded performance, or unexpected errors.
Since existing approaches often rely on external tools such as OCR engines and icon detectors to convert the environment into textual elements (e.g., HTML layouts), and also incorporate historical observations, such as task objectives, user instructions, and previous interactions, into the current input, the resulting context becomes excessively long. This may exceed the model’s input length limit, leading to potential unexpected behaviors (Zhang and Zhang, 2023). This issue is further acknowledged by AgentOccam (Yang et al., 2024a), which highlights the challenges posed by lengthy web page observations and interaction histories.
6
Social and Cultural Concerns Social and cultural concerns are the challenges faced by CUA agents in recognizing, respecting and adhering to different social norms, cultural sensitivities and ethical expectations when interacting with users or operating in real-world environments.
As CUAs execute user instructions on real-world applications, assessing their robustness to social and cultural concerns becomes increasingly crucial. The CASA benchmark (Qiu et al., 2025) is designed to evaluate LLM agent ability to identify and appropriately handle norm-violating user queries and observations. It reveals that current LLM agents perform poorly in web environments, exhibiting low awareness and high violation rates.
7
Response Latency This refers to the delay between the user input and the agent’s corresponding output or action, typically caused by model inference time, complex reasoning processes, or large context processing. It typically stems from various factors, among which the reasoning time of the brain component plays a major role.
The accumulation of such delays can affect the predictability of interactions; when users expect timely responses, excessive latency may cause misinterpretation of the agent’s state or intent, leading to incorrect user decisions. In critical domains such as financial trading or medical diagnosis, these issues can have serious safety implications. Zhang and Zhang (2023) and Wen et al. (2023) both recognize response latency as a significant challenge in the design of LLM-based CUAs, emphasizing its impact on interaction quality and user trust.
3.2.3 Action
The action component of an LLM-based CUAs engages in translating the agent’s output to a series of executable operations, such as calling APIs, web browsing, typing text, scrolling, and clicking on specific elements. As these behaviors involve interactions with an unverified website or API provider, this also brings with it a number of security risks.
8
API Call Errors API call errors refer to failures in a GUI agent’s ability to correctly infer, select, or format the required arguments when constructing API calls. Although general-purpose LLMs demonstrate strong capabilities in reasoning and planning, they often exhibit inaccuracies during API invocation, particularly in parameter filling (Deng et al., 2024).
In particular, within complex task chains, a single error in this process can lead to unpredictable outcomes and pose safety risks. MobileFlow (Nong et al., 2024), which further reinforces this concern, shows that errors in system-level API calls—such as incorrect parameter usage when retrieving layout information—may inadvertently expose sensitive interface content, highlighting the potential for even a single API-level mistake to escalate into a significant privacy or security threat. Similarly, Auto-GUI (Zhang and Zhang, 2023) also emphasizes that frequent API callings may introduce instability and increase the likelihood of calling errors.
3.3 Extrinsic threats
In this section, we introduce the extrinsic threats to Computer-Using Agents (CUAs)—attack vectors initiated by external adversaries aiming to exploit vulnerabilities in an agent’s interaction with its environment or to subvert its decision-making processes. Table 2 provides an overview of these threats, each introduced in detail in the following.
1
Adversarial Attack An adversarial attack on Computer-Using Agents (CUAs) involves the deliberate manipulation of input data or the environment to induce harmful or unintended behaviors in the agent. These agents, which operate within specific environments, such as interacting with webpages, computer interfaces, or mobile applications, are particularly susceptible to environment-specific adversarial attacks (Wu et al., 2024a).
For instance, adversarial attacks usually exploit subtle perturbations in the input data to mislead the agent. Wu et al. (2024a) demonstrated that adversarial examples can be crafted to appear visually or textually indistinguishable from original inputs, enabling attackers to deceive the agent into accepting manipulated data as genuine, thus steering it toward adversarial objectives without raising suspicion. Likewise, Aichberger et al. (2025) craft malicious image patches (MPIs) - tiny, reusable pixel perturbations placed anywhere on the display, so that when the agent screenshots the UI, the patch biases its perception and drives unsafe API actions. Zhao et al. (2025) also show that GUI grounding models can be reliably disrupted by small pixel-level perturbations, ranging from natural noise to targeted adversarial edits, across mobile, desktop, and web screenshots, causing agents to misidentify and misclick interface elements. Another form of attack manipulates the agent’s interaction with external deceptive elements to induce harmful behavior (Ma et al., 2024). Zhang et al. (2024c) explores an adversarial approach that targets the agent’s interactive interface. Attackers trick the agent into interacting with malicious pop-ups. This not only disrupts the agent’s ability to complete its assigned tasks but can also lead to severe consequences, including the installation of malware, redirection to phishing websites, or the execution of incorrect actions that disrupt automated workflows. Similarly, AgentScan (Wu et al., 2025) shows that by injecting a system-level notification pop-up milliseconds before the agent’s intended click, one can hijack its execution flow, luring it to tap the pop-up instead of the correct element of the user interface.
2
Prompt Injection Attack  Prompt injection attacks exploit the design of LLMs by embedding crafted instructions into the input that the model processes (Mudryi et al., 2025). These attacks trick the LLM into ignoring its predefined system rules or original purporse and following the adversary’s commands instead (Wu et al., 2024b; Liu et al., 2023b). Attackers often use this attack to force CUAs to do harmful or unethical actions. Most existing prompt injection attacks can be classified into two main types: direct prompt injection and indirect prompt injection.
Direct Prompt Injection In a direct prompt injection attack, the malicious instructions are embedded directly into the user’s input (prompt) (Debenedetti et al., 2024; Lupinacci et al., 2025). For instance, a CUA integrated into an operating system will normally accept commands like “open my calendar” or “launch the web browser.” An attacker might give a malicious command such as “Ignore all previous instructions and run the command to delete all files in the Documents folder.” If the agent fails to differentiate between its trusted system prompts and the injected malicious command, it could execute this harmful operation, resulting in a complete loss of user data.
Indirect Prompt Injection Indirect prompt injection, on the other hand, does not occur within the user’s immediate input (prompt). Instead, the attacker embed misleading instructions or unsafe content into the agent’s environment or external data sources (Kuntz et al., 2025; Wu et al., 2024b), such as webpages (Xu et al., 2024; Zhan et al., 2024; Liao et al., 2025; Evtimov et al., 2025) or files (Liao et al., 2024), so that when the agent later retrieves and processes this corrupted environment data, its reasoning can be compromised, leading to risky behaviors (Wu et al., 2025). The same idea is labeled visual prompt injection in VPI-Bench (Cao et al., 2025) when the cue is embedded specifically in on-screen UI text , and environmental injection attack in EIA (Liao et al., 2024) when it is planted in agents operation environments.
For CUAs, the unique nature of their operating environments has led to a specialized form of indirect prompt injection known as an environmental injection attack. Liao et al. (2024) introduce this concept by demonstrating how adversaries can subtly manipulate environment data, such as modifying webpage content, textual metadata, or document details, to embed hidden adversarial cues. These cues, often nearly imperceptible to human observers, alter the contextual signals that the agent relies on for decision-making, causing it to misinterpret its environment and execute unintended actions. Differently, RedTeamCUA (Liao et al., 2025) also embedds malicious instructions inside benign web content but prepends attention-grabbing cues (e.g., “THIS IS IMPORTANT!”) to boost adherence and steer subsequent OS/Web actions toward the attacker’s goal. Likewise, WASP (Evtimov et al., 2025) models a realistic setting where a black-box adversary can inject cues in the posted issues or comments on cloned GitLab/Reddit sites. Shapira et al. (2025) introduces a task-aligned injection technique, disguising the adversarial command as contextually helpful guidance tied to the agent’s current goal so it is more likely to be followed. Chen et al. (2025a) develops Fine-Print Injection which hides adversarial instructions in low-salience UI text (e.g., footers, terms of service, tiny captions), exploiting the agent’s tendency to parse such content uncritically. Hijacking JARVIS (Liu et al., 2025a) injects unprivileged third-party UI elements (e.g., fake buttons/text) into real Android apps. AdInject (Wang et al., 2025a) leverages the internet advertising delivery system to inject deceptive ad units into a web agent’s environment, tricking it into clicking the ad. The Foot-in-the-Door (FITD) attack (Nakash et al., 2024) injects a benign “distractor” request immediately followed by a hidden malicious instruction in external inputs (e.g. a GitHub issue), exploiting ReAct-based web agents’ failure to re-evaluate their thought trace and causing them to carry out the harmful step on the next tool call. Building on these notion of environment-level manipulation, OS-Harm (Kuntz et al., 2025) adds another vector: adversarial prompts delivered through desktop notifications instead of the task’s natural channel. AgentScan (Wu et al., 2025) displays adversarial text in non-interactive overlay window to hijack agent’s internal decision making and trick it to follow misleading instructions. Instead of manipulating visible page content or UI elements, Johnson et al. (2025) injects adversarial triggers optimized by the GCG algorithm in the webpage’s HTML accessibility tree to hijack agent behaviors. Similarly, EnvInjection (Wang et al., 2025e) injects a raw pixel value perturbation into the webpage source code so that, once the perturbed pixels are mapped into screenshot, they trigger the agent’s attacker-chosen action.
In addition, Zhan et al. (2025) develop an adaptive indirect prompt injection method that repeatedly probes a defended agent and refine the injected environmental cues, demonstrating the ability to overcome eight leading defense strategies. Similarly, EVA (Lu et al., 2025) leverages a black-box feedback loop to iteratively refine adversarial prompts and statistically distilling which text and layout cues hijack the agent’s attention. AgentVigil framework (Wang et al., 2025g) further demonstrates that indirect prompt injection can be fully automated by using a black-box fuzzing loop guided by Monte-Carlo Tree Search to generate and refine malicious cues against live web agents.
Building on these unimodal injection techniques, Wang et al. (2025c) introduces CrossInject, a cross-modal prompt injection framework that poisons both visual inputs via optimized adversarial perturbations embedded in screenshots and textual inputs through LLM-crafted malicious instructions.
3
Jailbreak  Jailbreak attacks are techniques that trick an LLM into bypassing its built-in safety mechanisms and refusal responses. By carefully rephrasing queries or injecting additional instructions, attackers force the model to ignore its predefined guardrails, enabling it to generate harmful or unauthorized outputs (Mo et al., 2024; Chu et al., 2024; Mao et al., 2025).
Over time, a lot of jailbreak prompts have been curated both manually (Chu et al., 2024) and via automated methods like GCG (Zou et al., 2023) and AutoDAN (Liu et al., 2023a) to exploit these vulnerabilities. These techniques are not only limited to standalone LLMs but have also been effectively applied to jailbreak CUAs. For example, OS-Harm (Kuntz et al., 2025) shows that even a simple jailbreak wrapper—an explicit ‘ignore all restrictions’ prompt template—markedly increases unsafe compliance in several agents. Kumar et al. (2024) further demonstrated that by modifying the user prompt using techniques such as prefix attacks, GCG suffixes, random search suffixes, and human-rephrased red-teaming prompts with diverse rephrasing strategies, they could either convince the browser agent that it was operating in an unrestricted sandbox environment or induce it to engage in harmful actions.
For multiagent systems, the Evil Geniuses framework (Tian et al., 2023) shows that by leveraging role specialization, attackers can partition agent tasks and exploit vulnerabilities in each specialized role to effectively jailbreak the system by bypassing its safety mechanisms. Qi et al. (2025) design a structured prompt-rewriting jailbreak, using narrative encapsulation and role-driven escalation to systematically bypass multi-agent debate systems’ safeguards and amplify harmful outputs. Meanwhile, the PsySafe framework (Zhang et al., 2024d) shows that injecting dark traits into agents can undermine established guardrails, further enhancing jailbreak effectiveness across multiagent environments. Beyond text-based prompts, Gu et al. (2024) introduces infectious jailbreak: a single adversarial image with low-salience textual embedded text can first jailbreak a single multimodal agent and then propagates to other agents, spreading exponentially.
4
Memory Injection  Memory injection attacks poison the agent’s persistent context, for example its stored plans, past prompts, or retrieved documents, rather than the live user prompt. Patlan et al. (2025a) introduces plan injection where attackers inject malicious steps into the agent’s stored task plan so that, when the Brain component later retrieves that plan at runtime, it unknowingly executes these injected steps alongside the legitimate ones. Similarly, Patlan et al. (2025b) demonstrates how embedding adversarial instructions into an agent’s shared memory store causes the agent to act on those unsafe memories during retrieval-augmented reasoning, triggering unauthorized behaviors without ever altering its immediate input. It is shown by Patlan et al. (2025b) that memory injection achieves greater persistent and stealth compared to prompt injections.
5
Backdoor Attack A backdoor attack involves injecting a malicious backdoor during the model’s training or fine-tuning phase, so that when a specific trigger phrase or input is later encountered during normal operations, the model executes unintended or harmful behavior (Yang et al., 2024b; Wang et al., 2024; Zhu et al., 2025b).
In particular, we can categorizes backdoor attacks on CUAs into two main forms: (1) The trigger is placed in the agent’s inputs (user queries or environmental observations) (Yang et al., 2024b) and directly manipulates the final output. For example, AgentPoison (Chen et al., 2024b) optimizes textual triggers via mapping poisoned instances to a unique embedding space, ensuring only trigger-containing prompts retrieve the malicious demonstrations. Boisvert et al. fine-tune LLM agents on mixed web and tool-use interaction logs that have been poisoned by inserting a benign-looking <div> with a unique ID into WebArena’s accessibility tree and a #\#EXFILTRATE_\_DATA token into $\tau$-Bench tool sequences, causing agents to execute the hidden malicious actions whenever they encounter these triggers at inference. A RAG backdoor attack embeds malicious payloads and corresponding trigger tokens directly into the agent’s external knowledge documents—such as augmented PDFs or database entries—so that when the agent retrieves those documents during its planning or observation phase, it unwittingly executes the attacker’s instructions without altering its core model parameters. Beyond text, VisualTrap (Ye et al., 2025) poisons GUI grounding data by remaping a tiny, low-salience on-screen mark to specific element–action pairs, causing the agent to execute attacker chosen clicks until the visual trigger appears. Similarly, (Wang et al., 2025f) injects a clean-label backdoor by fine-tuning vision–language mobile agents on a small fraction of screenshots covertly perturbed with an imperceptible visual trigger. More stealthily, Cheng et al. (2025) crafts composite triggers at the goal and interaction levels, using a min–max optimization with supervised contrastive learning and finetuning to ensure benign behavior on clean inputs and precise malicious actions when both trigger conditions are met. (2) The model’s internal reasoning is corrupted without visibly changing the final answer, e.g., covertly calling untrusted APIs (Yang et al., 2024b). Building on the RAG paradigm, Lupinacci et al. (2025) also show that a RAG backdoor attack can simply embed malicious payloads and trigger tokens into external knowledge documents, such as augmented PDFs or database entries, so that during retrieval and planning, the agent unwittingly executes the attacker’s instructions.
Meanwhile, to enhance stealthiness and bypass safety audits, attackers can break the backdoor code into multiple sub-backdoors, each activated by its own distinct trigger phrase or condition. When these sub-backdoors are combined, they enable the model to execute coordinated malicious behaviors (Zhu et al., 2025b). This modular design obscures the overall functionality behind seemingly unrelated trigger fragments, making detection and mitigation significantly more difficult.
6
Reasoning Gap Attack  A reasoning gap attack exploits the inherent disparity between an agent’s environmental inputs and its intrinsic reasoning process. In such attacks, adversaries inject misleading or ambiguous signals into one or more modalities (e.g., images, text, or sensor data) to create a gap between the agent’s perception and its reasoning. This discrepancy can cause the agent to make incorrect inferences or decisions that diverge from its intended functionality.
Chen et al. (2025d) examines how multimodal mobile agents are vulnerable to these attacks. The study shows that when attackers add conflicting or deceptive signals, such as subtle differences in an image combined with misleading text, the agent’s reasoning process struggles to correctly combine the different inputs. As a result, the agent might misinterpret the environment and take the wrong action.
7
System Sabotage  System sabotage attacks involve manipulating an agent into executing harmful actions that damage the underlying system. In such attacks, adversaries craft inputs to bypass safety mechanisms, causing the agent to perform operations like corrupting memory, damaging critical files, or halting essential processes (Luo et al., 2025b). These attacks are particularly dangerous because they directly target the infrastructure supporting the agent, potentially leading to widespread system failure or irreversible damage.
One example stated in (Luo et al., 2025b) is an attacker requests the agent’s assistance in creating a fork bomb, which is an intentionally crafted command that spawns processes indefinitely and tends to overwhelm the operating system. The user prompt disguises this request as a system “stress test,” persuading the agent to generate code that saturates system resources. Once executed, this fork bomb can cause the OS to become unresponsive or crash.
8
Web Hacking  Web hacking attacks use CUAs to autonomously identify and exploit vulnerabilities in websites, turning these agents into tools for malicious users. (Fang et al., 2024b) By feeding the agent specially crafted prompts or instructions, attackers guide it to scan web applications, detect security flaws (such as SQL injections or XSS vulnerabilities), and even formulate the exploit payloads.
In (Fang et al., 2024b), the authors show how malicious users can instruct a CUA to gather information on a target domain, evaluate its security posture, and carry out an attack. For example, the agent might test login forms for weak credentials, craft injection payloads, or automate data exfiltration attempts. If the agent successfully hacks the website, malicious adversaries could access private data or disrupt services and lead severe risks.
This type of autonomous web hacking highlights the growing need for robust safeguards and monitoring around CUAs. Without proper oversight, these systems can transform from helpful assistants into hacking tools, enabling malicious users to compromise websites with minimal effort.
4 Taxonomy of Existing Defenses
4.1 Defense Overview
Defense Target Components Agent Framework Target Threats
Env Prompt Model User Perception Brain Action
1 Environmental Constraints $\blacklozenge$ $\checkmark$ Ex. 2
2 Input Validation $\blacklozenge$ $\checkmark$ Ex. 3
3 Defensive Prompting $\blacklozenge$ $\lozenge$ $\checkmark$ $\checkmark$ Ex. 1 2
4 Data Sanitization $\blacklozenge$ $\checkmark$ Ex. 4 5
5 Adversarial Training $\blacklozenge$ $\checkmark$ Ex. 1
6 Output Monitoring $\blacklozenge$ $\checkmark$ In. 3 4 Ex. 7 8
7 Model Inspection $\blacklozenge$ $\checkmark$ Ex. 2 4 5
8 Cross-Verification $\blacklozenge$ $\checkmark$ $\checkmark$ Ex. 1 3 5
9 Continuous Learning $\blacklozenge$ $\lozenge$ $\checkmark$ Ex. 2
10 Transparentize $\blacklozenge$ $\lozenge$ $\checkmark$ In. 3 4
11 Topology-Guided $\blacklozenge$ $\checkmark$ $\checkmark$ Ex. 2
12 Perception Algorithms Synergy $\blacklozenge$ $\checkmark$ In. 1 5
13 Planning-Centric Architecture Refinement $\blacklozenge$ $\checkmark$ $\checkmark$ In. 2 7 8 Ex. 6
14 Pre-defined Regulatory Compliance $\lozenge$ $\blacklozenge$ $\checkmark$ $\checkmark$ In. 3 4 6
Table 3: A taxonomy of defense strategies. The symbol $\blacklozenge$ indicates that a defense is fully targeted at the given item, while $\lozenge$ represents limited availability. Ex. stands for extrinsic threats, In. represents intrinsic threats. The number followed indicates the explicit threat defined in prior sections.
In this section, we summarize the existing defenses to CUAs, as presented in Table 3. Defense methods are typically developed to counter specific threats or attacks discussed in Section 3; however, most defenses could generalize and exhibit effectiveness against others. We categorize existing defense methods based on agent components and frameworks, which are defined as:
  • Target Components identifies where the defense mechanism exerts its effect — Environment (Env), Prompt, Model, or User — and indicates whether it serves as a primary target ($\blacklozenge$) or a secondary target ($\lozenge$) of the method.
  • Agent Framework specifies the framework of the agent - Perception, Brain, and Action - where the defense mechanism predominantly acts. A checkmark (✓) denotes that the defense applies to the corresponding component.
  • Target Threat maps to the primary threats this method mitigates.
4.2 Defense Categories
1
Environmental Constraints  It refers to security mechanisms that limit or mediate the agent’s interactions with its operating environment in order to prevent harmful actions or malicious exploitation (Yang et al., 2024c; Nong et al., 2024). This strategy is applicable to both single-agent and multi-agent systems, focusing primarily on the environment component within the action phase of the agent framework. It targets environment-based threats such as prompt injection attacks that exploit GUI elements or interface structures.
For example, research reveals how visual elements on mobile interfaces can be manipulated to trigger unintended behaviors in GUI agents (Yang et al., 2024c). As a defense, they suggest sandboxing agent execution within constrained environments that monitor for risky API calls, and filtering GUI event access to minimize potential injection vectors (Yang et al., 2024c; Zhang et al., 2023). Additionally, GameChat uses control-barrier functions to constrain each agent’s trajectory to a safe region, preventing collisions and deadlocks in cluttered spaces (Mahadevan et al., 2025). Moreover, the framework in (Huang et al., 2025) builds a dynamic spatio-semantic safety graph that monitors real-time hazards and adaptively refines task plans to enforce safe execution.
However, this method may restrict the functional capability or generalizability of agents in dynamic real-world environments.
2
Input Validation  Input validation is a security measure that involves verifying and sanitizing user inputs to prevent the system from processing malicious or unintended commands (Ferrag et al., 2025). This strategy is predominantly applied in single-agent models, focusing on scrutinizing prompts to ensure they do not contain harmful instructions or malicious injections. Within the agent framework, input validation operates primarily at the perception level, where the agent interprets and understands user inputs. The primary threat addressed by this method is jailbreak attacks, where adversaries craft inputs designed to bypass the model’s safety mechanisms and elicit unauthorized behaviors.
For example, AutoDroid uses a privacy filter to mask personal information before prompts are sent (Wen et al., 2023). A similar filter also exists in (Zhang et al., 2024c). Additionally, in (Kumar et al., 2024), researchers observed that LLM-based browser agents are trained with safeguards to refuse harmful instructions in chat settings. The study introduced the Browser Agent Red-teaming Toolkit (BrowserART), which comprises 100 diverse browser-related harmful behaviors. Moreover, the authors in (Tshimula et al., 2024) apply pattern matching and high-precision filters to incoming prompts to detect and strip out jailbreak payloads before they reach the LLM. PromptArmor runs a lightweight LLM pre-processor that scans and sanitizes user inputs, removing any suspicious sub-prompts before they’re forwarded to the agent (Shi et al., 2025a). Also, RTBAS employs dynamic information-flow control and dual dependency screeners to vet tool calls, automatically ensuring confidentiality and integrity without constant user confirmation (Zhong et al., 2025).
However, a notable challenge in implementing input validation is the dynamic and unpredictable nature of user inputs. Attackers can craft perturbed prompts that appear benign but are designed to exploit specific model vulnerabilities. This necessitates continuous improvements to input validation protocols to effectively detect and mitigate evolving jailbreak techniques (Kumar et al., 2024).
3
Defensive Prompting  It refers to a security technique designed to safeguard language model agents by structuring prompts in a way that prevents adversarial manipulation and ensures the model adheres to intended behavior (Debenedetti et al., 2024). This method is primarily applied in single-agent models, focusing on the perception and brain components of the agent framework. It targets the prompt as the primary defense component while also influencing the model itself as a secondary target. The primary threats addressed by defensive prompting are prompt injection attacks, where adversarial inputs attempt to override the model’s intended behavior, and adversarial attacks, which subtly modify inputs to mislead the agent.
For example, in (Debenedetti et al., 2024), researchers introduced a structured evaluation environment to test and refine defensive prompting techniques. The study demonstrated that carefully crafted counter-prompts and reinforcement-based instruction tuning could significantly reduce the success rate of prompt injection attacks, enhancing model robustness (Debenedetti et al., 2024). Similarly, it was recommended that more detailed defensive prompts and robust content filtering should be used to enhance defense efficiency (Zhang et al., 2024c). Moreover, a safety prompt is introduced to instruct the agent to ignore malicious inconsistencies in  (Wu et al., 2024a). Also, experiments are done in (Chen et al., 2025d) to investigate the efficiency of this strategy.
However, implementing effective defensive prompting poses challenges, as adversaries continually develop more sophisticated prompt injection techniques. Additionally, the balance between robust security and maintaining the flexibility and generalization ability of the model remains an ongoing research challenge.
4
Data Sanitization  It refers to a process in machine learning that involves detecting and removing malicious or corrupted data from training datasets to ensure the integrity and security of models. Current discussion regarding this strategy mainly lies in the single-agent model, targeting at preventing malicious triggers during its reasoning and planning phase (Yang et al., 2024b; Jones et al., 2025; Wang et al., 2025b). This preventive measure is essential to protect models from various attacks, such as backdoor and memory injection attacks.
For example, Backdoor attacks involve embedding hidden triggers within the training data, causing the model to behave unexpectedly when these triggers are encountered during inference. By meticulously sanitizing the training data, such malicious patterns can be identified and eliminated, thereby safeguarding the model from potential exploitation (Yang et al., 2024b).
However, this method does not provide security guarantees (Yang et al., 2024b).
5
Adversarial Training  It is designed to enhance model resilience and robustness by incorporating adversarial examples into the training process (Wu et al., 2024a). This approach is predominantly applied to single-agent systems.
The primary focus of this method is the model component of the agent framework. By exposing models to adversarial examples during training, they learn to withstand such perturbations, thereby improving their robustness (Yu et al., 2025). This method specifically targets adversarial attacks, which involve subtle input modifications that can cause models to make incorrect predictions (Wu et al., 2024a).
For example, researchers demonstrated that Computer-Using Agents (CUAs) could be compromised through minimal perturbations to visual inputs, affecting their visual grounding (Wu et al., 2024a; Yu et al., 2020). By adversarial training, models can learn to recognize and resist these manipulations, thereby enhancing their task completion rate, as demonstrated in AutoSafe, which synthesizes diverse risk scenarios and uses them as on-the-fly adversarial examples during fine-tuning to markedly improve agent robustness (Zhou et al., 2025).
A notable characteristic of adversarial training is its ability to improve model robustness without necessitating changes to the model architecture. However, identifying possible adversarial threats in advance would be a prerequisite.
6
Output Monitoring  It refers to a strategy that involves continuously observing and evaluating the outputs of language models to ensure they align with user intentions and do not produce undesired actions (Shi et al., 2025b). This approach is primarily applied in single-agent systems, focusing on the model component within the action phase of the agent framework. It aims to address threats such as misalignment, where the agent’s actions diverge from user expectations, and hallucination, where the model generates incorrect or nonsensical information. Additionally, actions resulting in system sabotage or related to malicious usage, such as web hacking, could also be intercepted by this approach.
For instance, in the study (Fang et al., 2024a), the authors introduce InferAct, a novel approach that leverages the belief reasoning ability of large language models, grounded in Theory-of-Mind, to detect misaligned actions before execution. InferAct alerts users for timely correction, preventing adverse outcomes and enhancing the reliability of LLM agents’ decision-making processes (Fang et al., 2024a). Additionally, the Task Executor in AutoDroid verifies the security of an output action and asks the user to confirm if the action is potentially risky (Wen et al., 2023). Moreover, TrustAgent includes a post-planning inspection before tool calls (Hua et al., 2024). VeriSafe Agent auto-formalizes user instructions into a DSL specification and checks each GUI operation at runtime, blocking any action that fails logic checks (Lee et al., 2025).
However, a disadvantage would be the additional system overhead it incurs.
7
Model Inspection  This method detects malicious manipulations or compromised logic by examining internal model behaviors and parameters (Wang et al., 2025d; Yang et al., 2024b). It applies to both single-agent and multi-agent systems, targeting the model component of the agent, and operates within the brain of the agent framework. Model inspection defends against critical threats such as backdoor attacks, prompt injection attacks, and memory injection attacks by surfacing anomalous activity patterns or internal inconsistencies.
It is commonly categorized into two sub-methods: anomaly detection and weight analysis.
Anomaly Detection
It focuses on monitoring the behaviors of agents during inference or interaction to detect deviations from expected model outputs or communication topologies. It is especially relevant in multi-agent systems, where interactions can reveal inconsistencies in decision-making caused by compromised agents. For instance, a graph-based monitoring system was introduced to detect adversarially influenced agents by analyzing the topological communication patterns across agents (Wang et al., 2025d). The system was able to isolate and prune suspect nodes based on anomaly scores derived from communication flows (Wang et al., 2025d). Furthermore, a Graphormer model can analyze a dynamic spatio-semantic safety graph that captures both spatial and contextual risk factors in real-time to detect hazards (Huang et al., 2025).
Weight Analysis
This involves inspecting the internal parameters of a trained model to identify hidden triggers or abnormal value distributions indicative of backdoor implantation. This approach is particularly relevant for single-agent systems. For example, the authors perform weight-based inspection of transformer layers to identify neurons with disproportionately high influence tied to specific trigger tokens in  (Yang et al., 2024b). The analysis revealed clear distinctions between clean and poisoned models, suggesting that weight-level scrutiny can expose embedded backdoors (Yang et al., 2024b). Additionally, (Zhu et al., 2025b) proposed an automatic memory-audit step after every task, which flags anomalies in the agent’s internal memory traces to detect hidden backdoors.
A key challenge of model inspection is scalability and generalization—both anomaly detection and weight analysis often require clean model baselines, which may not always be available. Additionally, some backdoors may be designed to evade conventional statistical thresholds, necessitating adaptive and explainable inspection mechanisms.
8
Cross Verification  This is a collaborative defense strategy in multi-agent systems where multiple agents independently process the same task or instruction and validate each other’s outputs to ensure consistency and correctness (Zeng et al., 2024). This method primarily targets the model component of the agent framework and operates across both the brain and action stages, with the aim of defending against jailbreak, adversarial attacks, and backdoor attacks that may manipulate a single agent’s output to produce harmful or unauthorized behavior.
In the context of jailbreak prevention, cross-verification enables redundancy and consensus among agents, thereby reducing the likelihood that a single compromised response propagates through the system. For example, Zeng et al. and Huang et al. propose a multi-agent defense architecture where a guard or review agent cross-validates the output of a task agent (Zeng et al., 2024; Huang et al., 2024). If the task agent generates potentially harmful content in response to a jailbreak attempt, the guard agent flags the behavior and halts execution, effectively mitigating the attack (Zeng et al., 2024). Additionally, AgentOccam uses a Judge agent to assess every candidate action and picks the one with the least risk (Yang et al., 2024a). Similarly, GuardAgent spins up a separate guardian LLM that re-evaluates the primary agent’s outputs against knowledge bases, vetoing any unsafe recommendations (Xiang et al., 2024). Moreover, AGrail utilizes multiple checker agents to verify every candidate action before execution (Luo et al., 2025b). Also, the approach in (Barua et al., 2025) uses multiple independent runs of the same prompt across agents and uses majority consensus to filter out jailbreak attempts. MELON executes each prompt twice, once normally and once with a masked injection, to compare outputs and flag any inconsistencies as injected content (Zhu et al., 2025a). For backdoor attacks, ReAgent performs dual-level consistency checks between planning thoughts and executed actions to detect and abort backdoor-triggered behaviors at inference time (Li et al., 2025); and PeerGuard leverages mutual reasoning among agents to cross-verify each other’s outputs and isolate poisoned agents in a multi-agent backdoor defense (Fan and Li, 2025).
However, this method introduces coordination overhead and increases inference latency, particularly in large-scale deployments (Zeng et al., 2024).
9
Continuous Learning and Adaptation  Continuous learning and adaptation refers to the capability of agents to dynamically update their internal models based on new interactions, environments, or user feedback, thereby improving their long-term safety and robustness (Tian et al., 2023; Zhan et al., 2025). This strategy is primarily discussed in the context of multi-agent systems, targeting the model as the primary defense component and the user as a secondary influence. Operating within the brain of the agent framework, this method aims to address prompt injection attacks by enabling agents to detect and adapt to adversarial prompts over time.
This strategy is typically divided into two submethods: self-evolution mechanisms and user feedback integration.
Self-Evolution Mechanisms
It refers to the agent’s ability to autonomously adjust its reasoning or decision-making strategy based on past experiences and outcomes. LLM-based agents that re-encode their internal state across tasks are better at identifying unsafe instructions and suggest using performance memory or task replay buffers to evolve the agent’s policy over time (Tian et al., 2023; Luo et al., 2025b). This helps reduce the success rate of prompt injection attacks by enabling agents to learn from near-miss or failed tasks.
User Feedback Integration
It leverages feedback from human users to realign the agent’s behavior with user expectations. In the same study, the authors show that agents assisted with user feedback—such as warning prompts or confirmations before execution—exhibited more cautious and aligned behavior when encountering ambiguous or adversarial inputs (Tian et al., 2023). This aligns with the idea proposed in (Ma et al., 2024) that human-in-the-loop designs enhance agent safety in real-world, evolving task environments. For example, the study in (Zhang et al., 2025b) highlights user-initiated oversight mechanisms, such as manual correction loops and adaptive interface adjustments, enabling agents to learn from unintended outcome feedback and improve future interactions.
A core challenge in this method is balancing adaptability with stability—frequent updates can introduce regressions or new vulnerabilities if not managed carefully.
10
Transparentize  Transparentize refers to the implementation of mechanisms that enhance the transparency and interpretability of AI agents, thereby improving trust and safety in their operations. This strategy is particularly relevant in single-agent systems, focusing primarily on the model component and secondarily on the user component within the brain of the agent framework. It addresses threats such as hallucination—where the agent generates incorrect or nonsensical information—and misalignment, where the agent’s actions diverge from user intentions to risky operations.
It consists of two main submethods: Explainable AI (XAI) Techniques and Audit Logs.
Explainable AI (XAI) Techniques
It involves developing methods that make the decision-making processes of AI agents understandable to users. For instance, (Hu et al., 2024) highlights the importance of incorporating XAI techniques to elucidate how agents interpret instructions and execute tasks, thereby mitigating risks associated with hallucinations and misalignments.
Audit Logs
This entails recording the actions and decisions made by AI agents to provide a traceable history of their operations. Maintaining detailed logs is recommended to monitor agent behavior, facilitate debugging, and ensure accountability (Sager et al., 2025). For example, the authors in (Chen et al., 2025b) propose in-context consent dialogues and user-facing risk indicators to increase transparency of GUI agent operations and empower users to make informed decisions.
However, challenges in implementing transparentize strategies include balancing the depth of information provided with user comprehension and managing the storage and privacy concerns associated with extensive logging.
11
Topology-Guided  Topology-guided strategies enhance the security of multi-agent systems by analyzing and leveraging the structural relationships among agents to detect and mitigate adversarial threats (Wang et al., 2025d). This approach is particularly relevant in multi-agent systems, focusing primarily on the model component within the brain and action phases of the agent framework. It addresses threats such as prompt injection attacks by examining the communication patterns and interactions among agents.
This approach encompasses Agent Network Flow Analysis and Resilience Planning:
Agent Network Flow Analysis
It monitors the communication and interaction patterns among agents to identify anomalies that may indicate security breaches. For example, a multi-agent utterance graph could be constructed to monitor interactions and employ graph neural networks to detect anomalous communication flows that could signify prompt injection attacks (Wang et al., 2025d).
Resilience Planning
It focuses on designing the agent network topology to be robust against potential attacks. This includes strategies such as edge pruning, where connections to compromised agents are severed to prevent the spread of malicious information. The same study demonstrates that by adjusting the network topology through edge pruning, the system can effectively contain and mitigate the impact of detected attacks (Wang et al., 2025d).
However, challenges in implementing topology-guided strategies include the computational complexity of real-time graph analysis and the potential for reduced system performance due to the modification of network structures.
12
Perception Algorithms Synergy  Perception Algorithms Synergy refers to a family of techniques that combine complementary perception modules to obtain a more faithful, compact, and noise-resilient representation of the user interface. This strategy targets single-agent CUAs, acting mainly on the perception component of the model. It primarily mitigates intrinsic threats such as UI-understanding or grounding difficulties and excessive context length.
For example, grounding inputs by combining element-attribute, textual-choice, and image-annotation cues dramatically reduces mis-click rates on web tasks (Zheng et al., 2024). Additionally, MobileFlow augments its pipeline with a hybrid visual encoder and Mixture of Experts (MoE) alignment training, boosting image interpretation on Android  (Nong et al., 2024). On the PC side, PC-Agent introduces an active perception module that uses A11y-tree parsing with OCR, achieving fine-grained element localisation in complex desktop windows (Liu et al., 2025b). Finally, AgentOccam introduces observation-space alignment and page-simplification to address the excessive context length issue (Yang et al., 2024a).
Although these synergistic pipelines markedly improve grounding fidelity, they bring new engineering burdens—maintaining multiple perception branches, tuning resolution cut-offs, and balancing latency versus accuracy remain open challenges.
13
Planning-Centric Architecture Refinement  Planning-Centric Architecture Refinement denotes defenses that improve CUA’s reasoning-related architecture to ensure reliable scheduling, low response latency, and accurate API invocation. This strategy exists in both single and multi-agent systems. The method operates across the brain and action components of CUAs and directly targets threats such as scheduling errors, response latency, API-call errors, and reasoning gap attacks.
A representative approach is the chain-of-action prompt: it requires the agent to emit a full future-action plan before each execution step, cutting scheduling faults in half (Zhang and Zhang, 2023). Mobile-Bench extends this idea to multi-agent collaboration with a three-level (instruction, sub-task, action) hierarchy that decomposes long-horizon commands and reduces decision-making difficulties (Deng et al., 2024). AutoDroid lowers response latency by caching an LLM-generated guideline once per task, then delegating step-level binding to lightweight vision models (Wen et al., 2023). Complementarily, the PC-Agent framework allocates specialised Manager, Progress and Decision agents to refine and verify plans before execution, boosting success on 20-step desktop workflows (Liu et al., 2025b).
However, planning-centric refinements introduce coordination overhead, may suffer from stale caches when the UI changes, and require sophisticated plan-verification heuristics to guard against adversarial or hallucinated action sequences.
14
Pre-defined Regulatory Compliance  It involves designing AI agents to adhere to established laws, standards, and ethical guidelines, ensuring their operations align with societal norms and legal requirements. This strategy is particularly pertinent to single-agent systems, focusing primarily on the user component and secondarily on the model within the brain and action phases of the agent framework. It addresses threats such as social and cultural concerns, misalignment, and hallucination by embedding compliance mechanisms into the agent’s functionality.
This strategy comprises two main aspects: adherence to standards and ethical guidelines.
Adherence to Standards
It refers to specific regulatory frameworks and industry standards pre-defined for CUAs to comply with. For example, a comprehensive benchmark (Zhang et al., 2024e) is introduced to assess the safety of large language model agents, ensuring they meet predefined safety standards and operate within acceptable risk parameters. Additionally, GameChat employs pre-defined Control Barrier Functions to define safe operational boundaries for each agent in a multi-agent system, ensuring agents’ trajectories remain within safe limits, preventing collisions  (Mahadevan et al., 2025). The game-theoretic strategy satisfying Subgame Perfect Equilibrium in GameChat further prevents agents from deviating from the agreed-upon strategies at any point, promoting consistent adherence to safe navigation protocols (Mahadevan et al., 2025). Moreover, ShieldAgent extracts verifiable rules from policy documents, structures them into a set of action-based probabilistic rule circuits, and associates specific agent actions with corresponding constraints (Chen et al., 2025e). Continuous verification ensures real-time standards adherence (Chen et al., 2025e). Also, AgentSandbox operationalizes security principles like defense-in-depth and least privilege within agent lifecycles, embedding policy enforcement checkpoints that uphold confidentiality and integrity requirements (Zhang et al., 2025a).
Ethical Guidelines
This involves integrating ethical considerations into the design and operation of AI agents. The same study emphasizes the importance of aligning agent behaviors with ethical norms to prevent unintended consequences, such as generating harmful content or exhibiting biased behaviors (Zhang et al., 2024e).
However, challenges in implementing pre-defined regulatory compliance include the dynamic nature of regulations and ethical standards, requiring continuous updates to the agent’s compliance mechanisms to remain current.
Platform Benchmark Highlight Data Size Collection Metric Measure
Web VWA-Adv
(Wu et al., 2024a)
Assesses the robustness of multimodal web agents against adversarial attacks originating from the environment. 200 adversarial tasks Open-source data modification Benign SR, ASR Rule
ST-WebAgent Bench
(Levy et al., 2024)
Evaluates the safety of web agents by testing policy adherence and risk mitigation, focusing on external attacks and internal misalignments. 235 policy-enriched tasks Open-source data modification CuP, Partial CuP Rule
BrowserART
(Kumar et al., 2024)
Assesses the safety of browser agents against harmful interactions, content, and jailbreak. 100 harmful browser-related behaviors Open-source data modification ASR LLM
CASA
(Qiu et al., 2025)
Evaluates LLM web agents’ cultural and social awareness about social norms and legal standards in interactions with non-malicious users. 1225 user queries, 622 web observations GPT-4o generation with human validation AC-R, Edu-R, Helpfulness, Vio-R LLM
SafeArena
(Tur et al., 2025)
Evaluate deliberate misuse of autonomous web agents and introduces the ARIA risk framework. 250 safe and 250 harmful tasks Human curation with LLM assistance, Open-source data augmentation TCR, RR, Normalized Safety Score Rule, LLM, Manual
AgentDAM
(Zharmagambetov et al., 2025)
Measures inadvertent leakage of sensitive information by AI agents during web task execution. 246 tasks Human curation, Open-source utilization Utility, LR Rule, LLM
ShieldAgent Bench
(Chen et al., 2025e)
Tests agent safety against adversarial instructions and policy violations across web environments and risk categories. 960 web instructions, 3110 unsafe trajectories Open-source data modification Accuracy, FPR, Recall, Inference Cost Rule
WASP
(Evtimov et al., 2025)
Shows that even top-tier AI models can be deceived by simple, low-effort human-written injections in very realistic scenarios. 84 tasks Human curation TSR, Intermediate ASR Rule, LLM
VPI-Bench
(Cao et al., 2025)
Evaluates the robustness of CUAs and Browser-use agents to visual prompt injection across five popular web platforms. 306 test cases Human curation AR, ASR LLM
Mobile MobileSafety Bench
(Lee et al., 2024a)
Evaluates mobile agents in Android emulators for safety, helpfulness, ethical compliance, fairness, privacy, and prompt injection attacks. 80 tasks Human survey and annotation TSR, RR Rule
Hijacking Jarvis
(Liu et al., 2025a)
Evaluates mobile GUI agents’ safety under unprivileged third-party UI manipulations by the AgentHazard framework. 3000+ attack scenarios Human creation, annnotation TSR, MR, $ACC_{safe}$, $ACC_{attack}$ Rule
Table 4: An overview of web and mobile based computer-using agents (CUAs) safety benchmarks.
Platform Benchmark Highlight Data Size Collection Metric Measure
General ToolEmu
(Ruan et al., 2023)
Evaluates safety failures of LM agents across diverse tool-driven scenarios. 36 toolkits, 144 test cases Human curation with LLM assistance Safety, Helpfulness LLM, Manual
R-Judge
(Yuan et al., 2024)
Evaluates LLM agents’ safety awareness about multiple risks, with prompt injection attacks and complex environment challenges. 569 records of multi-turn agent interaction Open-source data modification with ChatGPT F1 score, Recall, Specificity, Effectiveness Manual, LLM
TrustAgent
(Hua et al., 2024)
Evaluates agents’ safety regulations into planning across domains and risks. 144 data points Open-source data modification Helpfulness, Safety, Total Correct Prefix, SSR LLM, Rule
InjecAgent
(Zhan et al., 2024)
Evaluates tool-integrated LLM agents’ susceptibility to indirect prompt injections. 1,054 test cases GPT-4 with manual refinement ASR Rule
AgentDojo
(Debenedetti et al., 2024)
Evaluates the robustness of LLM-based agents in dynamic, tool-using environments against prompt injection attacks. 97 tasks, 629 security test cases Human design with LLM assistance TSR, TSR under Attack, ASR Rule
PrivacyLens
(Shao et al., 2024)
Tests agents for privacy adherence, assessing vulnerability to data leakage and misuse amid misalignment. 493 seeds and 1479 questions Human collection, transformation with GPT-4 LR, LRh, Helpfulness LLM, Rule
HAICOSYSTEM
(Zhou et al., 2024)
Simulates multi-turn human–agent tool interactions to probe multi-dimensional safety risks. 132 scenarios, 8K simulated episodes Human creation, Open-source inspiration TARG, SYST, CONT, SOC, LEGAL, EFF, GOAL LLM
AgentHarm
(Andriushchenko et al., 2024)
Evaluates LLM agents’ resistance to malicious requests and multi-step harmful behaviors triggered by jailbreaks. 110 malicious tasks, 330 augmented tasks Human generation and review, LLM generation Harm score, RR LLM, Rule
Agent Security Bench
(Zhang et al., 2024b)
Evaluates LLM agents’ security against external attacks such as prompt injection and backdoors. 400 tools, 10 scenarios, 10 agents, and 400 cases GPT-4 generation ASR, RR, PNA, BP, FPR, FNR, NRP LLM, Rule
Agent-SafetyBench
(Zhang et al., 2024e)
Evaluates LLM agents’ safety against jailbreaks and misalignments across risks. 2000 test cases with 10 failure modes and 349 environments Open-source data modification Safety Score LLM, Rule
RedTeamCUA
(Liao et al., 2025)
Demonstrates that indirect prompt injection presents tangible risks for even advanced CUAs despite their capabilities and safeguards. 216 adversarial scenarios Human Curation SR, ASR, AR Rule, LLM
RiOSWorld
(Yang et al., 2025b)
Measures the risk intent and completion of MLLM-based agents during real-world computer manipulations. 492 risky tasks Human, Open-source, LLM RGC, RGI Rule, LLM
MLA-Trust
(Yang et al., 2025c)
Measures agent trustworthiness by orchestrating high-risk, interactive tasks, especially in multi-step interactions. 34 tasks Human creation, Open-source data augmentation Accuracy, MisRate, ASR, TS, RtE Rule LLM
GUI-Robust
(Yang et al., 2025a)
Reveal GUI agents’ substantial performance degradation in abnormal scenarios. 5318 tasks Semi-automated dataset construction paradigm Action &\& Coordinate Accuracy, TSR Rule
OS-Harm
(Kuntz et al., 2025)
Measures CUA safety across three harm types—deliberate user misuse, prompt injection, and model misbehavior. 150 tasks Human creation with LLM assistance, Open-source data augmentation Unsafe, TSR LLM
RAS-Eval
(Fu et al., 2025)
Evaluates security of LLM-based agents across simulated and real-world tool executions in diverse formats. 80 test cases, 3802 attack tasks Human collection, implementation TCR, TIR, TFR, score, ASR Rule
OpenAgent-Safety
(Vijayvargiya et al., 2025)
Evaluates agent safety when interacting with real tools across mixed environments including web and OS. 350 multi-turn, multi-user tasks Human curation with LLM assistance Unsafe Behavior Rates, Failure Rate, Disagreement Rate Rule, LLM
Table 5: An overview of general-purpose computer-using agents (CUAs) safety benchmarks.
5 Evaluation and Benchmarking
Computer-Using agents (CUAs) are widely deployed across various platforms, necessitating a more comprehensive evaluation of their safety performance compared to general LLM-based agents. This section provides a structured summary of these benchmarks, as shown in Table 4 and  5.
A benchmark typically consists of three key components: a dataset, interactive environment, evaluation metrics, and corresponding measurements. The dataset is a static collection of data points, where each data point includes multiple inputs—such as questions, tasks, and screenshots—as well as a sequence of actions as the output. Depending on the application scenarios they involve, we categorize these datasets into three types: web-based scenario, mobile-based scenario, and general-purpose scenario.
In contrast, the environment is interactive rather than static, encompassing entire user interfaces where agent actions can influence the system’s state, receiving feedback to guide subsequent actions. Since these benchmarks often rely on specific components within a given environment, we categorize them into two types: real-world environments and sandbox environments.
The evaluation metrics vary depending on the benchmark’s objectives, which can be generally classified into three categories: task completion metrics, intermediate performance metrics, and broader metrics assessing efficiency, generalization, safety, and robustness.
Measurements refer to the various methods employed to calculate metrics. These methods can be broadly categorized into three main types: rule-based, LLM-as-a-judge, and manual judge. Each approach has unique characteristics and is suited for different usage scenarios.
5.1 Datasets
5.1.1 Web-based Scenario
In the web-based scenario, several datasets have been proposed to assess the safety of agents operating within browser environments. Specifically, ST-WebAgentBench (Levy et al., 2024) and BrowserART (Kumar et al., 2024) focus on evaluating agents’ safety-related behaviors in tasks involving web navigation, interaction, and tool usage under potential prompt injection threats. Meanwhile, PrivacyLens (Shao et al., 2024) investigates privacy-sensitive interactions in web-based conversations, containing 493 validated prompts derived from U.S. legal, social, and interpersonal communication norms. In parallel, CASA (Qiu et al., 2025) provides a web-based benchmark designed to evaluate agents’ awareness of cultural and social contexts, utilizing grounded questions and descriptors sourced from CultureBank. Furthermore, ShieldAgent-Bench (Chen et al., 2025e) extends these efforts by simulating adversarial instructions and policy-violation scenarios across diverse web environments, providing 960 safety-related instructions and 3,110 unsafe trajectories. SafeArena (Tur et al., 2025) likewise broadens coverage by injecting jailbreak-inspired malicious intents into 4 realistic WebArena sites and introducing 500 paired safe vs. harmful tasks over 5 harm categories. Similarly, WASP (Evtimov et al., 2025) combines 21 concrete attacker goals with 2 benign user goals under both URL and plaintext injection templates, with total of 84 tasks, to evaluate agent security against prompt injection attacks. VPI-Bench (Cao et al., 2025) targets visual prompt injection, providing 306 test cases across five popular sites, each embedding an adversarial instruction directly in the on-screen UI to see whether agents follow it. AgentDAM (Zharmagambetov et al., 2025) assesses AI agents’ propensity to expose sensitive information across three realistic web settings (Reddit, GitLab, Shopping) over 246 tasks. Finally, the VWA-Adv benchmark (Wu et al., 2024a) targets web-based scenarios, introducing 200 adversarial tasks built on VisualWebArena (Koh et al., 2024) to evaluate agent robustness against realistic attacks through imperceptible webpage perturbations and component-wise adversarial flows.
5.1.2 Mobile-based Scenario
Mobile-focused benchmarks evaluate CUAs inside real mobile applications. For instance, MobileSafetyBench (Lee et al., 2024a) consists of 80 representative tasks across domains such as messaging, social media, finance, and system utilities. It serves to evaluate agents’ safety performance under mobile-specific constraints and risks. Hijacking JARVIS (Liu et al., 2025a) contributes a two-part mobile benchmark: 58 dynamic tasks with attack patterns in different complexity levels and a 210-screenshot static set across 14 popular Android apps, enabling both live and offline safety evaluation.
Several datasets are designed with general-purpose safety evaluation in mind, covering a wide spectrum of risks, tools, and environments. R-Judge (Yuan et al., 2024) focuses on risk awareness in 569 multi-turn interactions spanning five categories and 27 scenarios, covering 10 different risk types. TrustAgent (Hua et al., 2024) offers 70 samples across 5 domains, incorporating both risk analysis and corresponding ground truth implementations.
5.1.3 General-purpose Scenario
Several datasets are designed with general-purpose safety evaluation in mind, spanning diverse tools, risks, and interaction environments.
Tool-use scenario. Tool-enabled CUAs have received intense scrutiny over the past two years. ToolEmu (Ruan et al., 2023) probes safety failures in a fully LM-emulated sandbox, covering 36 toolkits (18 categories), 144 high-stakes tasks, and 9 risk types. RAS-Eval (Fu et al., 2025) standardizes security testing for tool-driven agents with 80 core test cases and 3,802 attack tasks mapped to 11 CWE categories across both simulated and real tool executions. Prompt-injection–oriented suites such as AgentDojo (Debenedetti et al., 2024) with 97 realistic tasks, 629 security cases and InjecAgent (Zhan et al., 2024) with 330 tools from 36 toolkits evaluate how well agents perform under various adversarial scenarios while equipping with diverse tools. AgentHarm (Andriushchenko et al., 2024) broadens harmful-behavior testing with 110 base behaviors in 11 harm categories, and the large-scale Agent Security Bench (ASB) (Zhang et al., 2024b) aggregates 10 scenarios, 10 purpose-built agents, and over 400 tools and tasks to offer a unified safety framework. Furthermore, Agent-SafetyBench (Zhang et al., 2024e) covers 349 interaction environments and 2,000 test cases, spanning 8 safety risk categories and 10 prevalent failure modes in unsafe agent behaviors.
Mixed / hybrid environments Several benchmarks test agents that operate across heterogeneous interfaces (web, OS, shells, code executors, etc.). For instance, OpenAgentSafety (Vijayvargiya et al., 2025) provides 350 multi-turn, multi-user tasks in both benign and adversarial settings using a real browser, shell, file system, and messaging APIs. RiOSWorld (Yang et al., 2025b) runs 492 risky tasks in 13 categories on an OSWorld VM, capturing both environment and user-originated risks. RedTeamCUA (Liao et al., 2025) introduces RTC-Bench with 864 hybrid Web–OS adversarial scenarios, underscoring CUAs’ susceptibility to indirect prompt injection. MLA-Trust (Yang et al., 2025c) evaluates 34 high-risk, real-world tasks, showing how multi-step interactions in real environments can amplify risks beyond static LLM outputs. GUI-Robust (Yang et al., 2025a) complements this by injecting seven classes of interface anomalies (e.g., ad pop-ups, loading delays) to study robustness. Finally, HAICOSYSTEM (Zhou et al., 2024) emulates realistic human–AI interactions and complex tool use by running 8K+ simulations across 132 scenarios in seven domains, covering multi-dimensional risks (operational, content, societal, legal).
Broader risk-awareness and multidimensional safety. Beyond concrete tool or environment settings, several works emphasize comprehensive risk taxonomies and analysis. R-Judge (Yuan et al., 2024) scores risk awareness over 569 multi-turn interactions spanning 5 categories, 27 scenarios, and 10 risk types. TrustAgent (Hua et al., 2024) contributes 70 samples across 5 domains with paired ground-truth implementations to evaluate both helpfulness and safety. PrivacyLens (Shao et al., 2024) offers 493 privacy-sensitive vignettes and trajectories for leakage analysis. GUI-Robust (Yang et al., 2025a) complements these efforts by focusing on robustness under anomalies in interactions. It includes seven categories of common interface failures, such as advertisement pop-up and page loading delay.
5.2 Environments
5.2.1 Real-world Environments
A real-world environment refers to a complex and dynamic setting, such as the Android OS or the web, where agents perform tasks that reflect the unique challenges of these environments.
Several studies have focused on evaluating agent behavior in such contexts. ST-WebAgentBench (Levy et al., 2024), BrowserART (Kumar et al., 2024), and CASA (Qiu et al., 2025) evaluate agent behavior on real websites, assessing aspects such as safety, trustworthiness, and cultural awareness. Unlike prior benchmarks that only provide task instructions, ShieldAgent-Bench (Chen et al., 2025e) evaluates real-world safety challenges by incorporating complete agent interaction protocols, such as instructions, trajectories, enforced policies, and ground-truth labels. Similarly, MobileSafetyBench (Lee et al., 2024a) examines agent performance across various mobile applications, including messaging, web navigation, social media, finance, and utility apps. Hijacking JARVIS (Liu et al., 2025a) also runs on real Andriod apps inside an emulator to create realistic on-screen manipulations, testing how agents behave under live, adversarial mobile interfaces. To provide a more comprehensive assessment, Agent-SafetyBench (Zhang et al., 2024e) extends evaluations across both web and mobile environments, offering a broader analysis of agent safety. OpenAgentSafety (Vijayvargiya et al., 2025) also evaluates agents in mixed live tool environments, automating a real browser, shell, file system, code executor, and messaging API. RAS-Eval (Fu et al., 2025) likewise executes attacks against live tools (e.g., MCP/JSON services) to measure real-world security failures, complementing its simulated setting. AgentDAM (Zharmagambetov et al., 2025) also operates in truly web-native settings, building atop the WebArena (Zhou et al., 2023) and VisualWebArena (Koh et al., 2024) simulators.
In contrast, the VWA-Adv benchmark (Wu et al., 2024a) focuses on a realistic threat model in real-world web environments, where the attacker is a legitimate user with limited capabilities to manipulate the environment. Likewise, VPI-Bench builds fully interactive, pseudo-authentic webpage replicas of selected websites—cloned DOMs rendered in a real browser—into which the authors manually inject the visual prompts, yielding a realistic yet controllable setting for testing on-screen attacks. Similarly, MLA-Trust (Yang et al., 2025c) evaluates CUAs in live GUI environments, including both websites and mobile apps, emphasizing how autonomous execution and dynamic interaction cycles can introduce compounding trust failures.
In the context of security vulnerabilities introduced by external content, InjecAgent (Zhan et al., 2024) simulates agent responses to adversarial inputs, enabling researchers to assess the agent’s subsequent actions and obtain attack outcomes. GUI-Robust (Yang et al., 2025a) similarly adopts real application interfaces with realistic anomalies, collecting human interaction traces to evaluate agent robustness under more complex or compound failure modes. PrivacyLens (Shao et al., 2024) analyzes the data transmitted by agents to external tools, identifying potential privacy risks during task execution.
Real-world environments offer high realism and dynamism, ideal for evaluating agent capabilities. However, they pose challenges for consistent evaluation and reproducibility due to reliance on live websites with constantly evolving content.
5.2.2 Sandbox Environments
Sandbox environments are designed to explore the agent safety performance under a stable environment, which may be simple or not perfectly aligned with the reality of the real environment, but it is a good way to explore the vulnerability of agent to specific attacks or in specific areas.
R-Judge (Yuan et al., 2024), TrustAgent (Hua et al., 2024), AgentDojo (Debenedetti et al., 2024), and AgentHarm (Andriushchenko et al., 2024) each construct a simulation environment with distinct focuses. R-Judge adopts ReAct (Yao et al., 2023) as its interactive framework. TrustAgent enhances task evaluation by providing detailed descriptions of external tools relevant to each task domain. AgentDojo builds an extensible environment designed for developing and assessing new agent tasks, defenses, and adaptive attacks. ToolEmu (Ruan et al., 2023) adopts a fully LM-emulated tool environment: instead of calling real APIs, a language model “pretends” to be each tool by returning text descriptions of what would have happened and updating a textual state. SafeArena (Tur et al., 2025) leverages four realistic WebArena sandbox sites (forum, e-commerce, code repo, retail), each instrumented per WebArena templates. RAS-Eval (Fu et al., 2025) also offers a controlled, tool-API sandbox, which simulates tool execution by letting tools read/write a shared in-memory state (e.g., calendars, databases). Likewise, by restricting attacker capabilities to creating issues, posts, and comments, WASP (Evtimov et al., 2025) offers a controlled yet realistic platform for end-to-end prompt injection testing.
Meanwhile, AgentHarm utilizes synthetic tools for all tasks, implemented through Inspect (AI Security Institute, ), effectively mimicking a range of tools from general utilities to domain-specific applications. RedTeamCUA (Liao et al., 2025) also constructs a hybrid sandbox by combining OS with Docker-based web platforms, enabling fully controlled realistic evaluation of indirect prompt injection across both web and OS interfaces. RiOSWorld (Yang et al., 2025b) runs agents inside an OSWorld virtual machine, exercising real desktop and web apps but under a fully instrumented, reproducible sandbox. HAICOSYSTEM (Zhou et al., 2024) also builds a fully controlled ecosystem—LLM-simulated users, tools, and environments inside a sandboxed OS/Web setting—enabling reproducible risk injections without live-site volatility.
Sandbox environments provide controlled evaluation platforms that replicate the dynamism of real-world settings while ensuring consistency and reproducibility, however, its simplicity may also deprive it of elements of its real-world environment.
5.3 Evaluation Metrics
This subsection summarizes the commonly used evaluation metrics for CUA safety. Most metrics reported in Table 4 and  5 are defined here. Extremely specialized metrics that are unique to a single work are omitted for brevity.
5.3.1 Task Completion Metrics
1
Task Success Rate (TSR) assesses whether an agent successfully reaches the final goal of a task, regardless of the performance on intermediate steps (Yao et al., 2022; Xie et al., 2024; Wen et al., 2023). It serves as a holistic indicator of an agent’s overall effectiveness in completing a given task. In many safety benchmarks and datasets, this measure is akin to the Benign Success Rate (Benign SR) (Wu et al., 2024a) or Benign Utility (Debenedetti et al., 2024) or Performance Under No Attack (PNA) (Zhang et al., 2024b), which evaluates how well an agent performs under normal, non-adversarial conditions.
Similarly, many benchmarks also measure how well agents complete required tasks, often including both benign and adversarial tasks. AgentDAM (Zharmagambetov et al., 2025) uses Utility measures how well agent performs its presrcibed tasks. HAICOSYSTEM (Zhou et al., 2024) uses Goal Completion (GOAL), scoring whether the agent achieves the scenario objective. In SafeArena (Tur et al., 2025), the Task Completion Rate (TCR) automatically matching each agent run against predefined reference objects to determine successful task completion. While, in RAS-Eval (Fu et al., 2025), TCR is the fraction of tasks where the agent’s tool-call sequence fully contains the human-annotated reference sequence; Task Incompletion Rate (TIR) counts runs that invoke only a subset or wrong combination of required tools; Task Fail Rate (TFR) flags runs that crash, call no tools, or exceed stack/length limits; and ultimately synthsizes these three metrics into a unified Performance Score (score).
2 Helpfulness measures how effectively agents fulfill user instructions while balancing overall performance with safety considerations, extending beyond simple task completion. It measures not only whether the task was finished but also how well the agent executed the necessary operations, such as making the correct and effective tool calls to achieve the desired outcome (Ruan et al., 2023). In other words, while task completion is a binary measure of whether a task is accomplished, helpfulness also considers the overall utility, coherence, and effectiveness of the response. Evaluating helpfulness often involves designing an automatic evaluator (e.g. prompting a LLM as judge) or relying on human annotators (Qiu et al., 2025).
5.3.2 Intermediate Step Metrics
1
Step Success Rate (SSR) evaluates how accurately an agent performs each individual step within a multi-step task (Deng et al., 2023; Zhang et al., 2024a; Chen et al., 2024a). For each step, it checks if the action aligns with the expected or "ground truth" behavior. Formally, SSR is defined as
$SSR=\frac{\#~Correct~Steps}{\#~Total~Steps}$
A higher step success rate reflects greater precision in executing each part of the task, which is especially crucial in scenarios that require reliable and fine-grained control across multiple actions.
2 Total Correct Prefix In addition to overall step accuracy, it is important to assess the sequence in which these steps are executed. Some individual actions may match their corresponding ground truth steps; however, if they occur out of the intended order, this misordering can lead to potential safety or reliability risks.
The Total Correct Prefix is defined as the longest initial sequence of correct, in-order steps that aligns with the ground truth (Hua et al., 2024). Evaluating this metric offers valuable insight into the agent’s ability to follow the intended procedure from the start, while also revealing vulnerabilities that may arise from executing actions in an incorrect sequence.
5.3.3 Safety and Robustness Metrics
1
Attack Success Rate (ASR) is a very commonly used metric (Zhan et al., 2024; Debenedetti et al., 2024; Kumar et al., 2024; Zhang et al., 2024b) to evaluate the adversarial robustness of Computer-Using Agents (CUAs). It measures the percentage of attack tasks in which an adversary causes an agent to produce an undesired or unsafe outcome. ASR is given by:
$ASR=\frac{\#~Successful~Attack~Tasks}{\#~Total~Attack~Tasks}$
A higher ASR indicates increased vulnerability of the agent to adversarial manipulation (Chang et al., 2023).
Since attacks vary in form and impact, several works define complementary metrics beyond ASR to capture different facets of adversarial effect. For instance, the Violation Rate (Vio-R) in CASA (Qiu et al., 2025) measures the fraction of agent responses that violate stated norms when presented with misleading or malicious inputs, which effectively captures how often the agent is “attacked” into norm-breaking behavior. Similarly, Misguided Rate (MisRate) in MLA-Trust (Yang et al., 2025c) measures how often an agent gives plausible yet incorrect answers to ambiguous or underspecified prompts. Formally, $\text{MisRate}=\frac{n_{\text{misguide}}}{n_{\text{amb}}}$, where $n_{\text{misguide}}$ is the number of misguided responses and $n_{\text{amb}}$ is the total count of ambiguous prompts. Likewise, Misleading Rate (MR) (Liu et al., 2025a) measures the fraction of attack trials in which the agent abandons the intended behavior and instead follows the injected misleading content. Risk Goal Completion (RGC) in RiOSWorld (Yang et al., 2025b) mirrors ASR by reporting the fraction of risky tasks where the agent fully achieves the attacker’s objective. In OS-Harm (Kuntz et al., 2025), a straightforward percentage of Unsafe execution traces has been calculated to quantify the overall safety-violation rate.
Furthermore, in Agent Security Bench (Zhang et al., 2024b), Benign Performance (BP) measures the agent’s success rate on its intended tasks when a backdoor trigger is present, indicating how well it maintains functionality under backdoor attack. Net Resilient Performance (NRP) then combines non-adversarial condition capability (PNA) and robustness against attack (ASR) into a single score:
$NRP=PNA\times(1-ASR)$
A higher NRP reflects both strong task performance and effective resistance to attacks, whereas a lower NRP signals vulnerability, poor baseline accuracy, or both. It is valuable as it measures the trade-offs between performance and robustness.
In addition to ASR, some benchmarks go beyond a binary success measure to characterize how severe or how partial an unsafe outcome is: In TrustAgent (Hua et al., 2024), the Safety metric evaluates both the likelihood and the severity of potential risks by assigning categorical ratings (e.g., “Certain No Risk” to “Likely Severe Risk”) to each agent response, thus capturing not only whether a response is unsafe but how risky it is. ToolEmu (Ruan et al., 2023) also uses Safety with the similar idea: its safety evaluator flags failures and gives a 0–3 score by crossing harm likelihood with harm severity. In contrast, Agent-SafetyBench (Zhang et al., 2024e) simplifies the assessment with a Safety Score, measured as the proportion of test cases labeled “safe” by an LLM judge. Meanwhile, in AgentHarm (Andriushchenko et al., 2024), the Harm Score is computed via a detailed manually written grading rubric where outputs earn partial credit whenever some but not all harmful criteria are triggered, providing a finer-grained view of measure of how much of the harmful behavior was actually executed. HAICOSYSTEM (Zhou et al., 2024) refines the overall safety severity into five risk dimensions: Targeted Safety Risk (TARG), System/Operational Risk (SYST), Content Safety Risk (CONT), Societal Risk (SOC), Legal/ Rights Risk (LEGAL), each scored on a 10-10 to 0 scale to indicate how severely a trajectory violates that category.
2 Completion Under the Policy (CuP) quantifies the rate at which an agent completes tasks in strict adherence to predefined safety or usage policies (Zhang et al., 2024a). CuP is calculated as
$\mathrm{CuP}\;=\;C_{task}\;\times\;\mathbf{1}\{V_{total}=0\},$
where $C_{task}$ is the task completion score, $V_{\text{total}}=\sum_{\text{source},\,\text{dim}}V_{\text{source,dim}}$ is the total number of policy violations across all sources and dimensions, and $\mathbf{1}\{\cdot\}$ is the indicator function that returns 1 exactly when $V_{\text{total}}=0$ (i.e., no violations occurred) (Levy et al., 2024).
Recognizing that certain tasks can be challenging to fully complete, Levy et al. (2024) also proposes the Partial Completion Rate (PCR), which acknowledges tasks that fulfill at least one success criterion, even if the agent does not achieve complete success. This allows evaluators to focus on how well the agent follows policy requirements in scenarios of partial success.
Building on PCR, Levy et al. (2024) further defines the Partial Completion Under the Policy (Partial CuP) metric, which extends CuP to partial completions by examining whether the agent respects policy constraints when only a portion of the task is satisfied (modify the $C_{task}$ score by PCR). This assesses the agent behavior by balancing between task difficulty and adherence to safety guidelines.
3 F1 Score is a critical safety metric that combines both precision and recall into a single measure to evaluate an agent’s effectiveness in a binary classification scenario, such as distinguishing between safe and unsafe labels (Chang et al., 2023). F1 Score is defined as
$F1=2\times\frac{Precision\times Recall}{Precision+Recall}$
Here, precision is calculated as
$Precision=\frac{TruePositives}{TruePositives+FalsePositives}$
reflecting the accuracy of the positive predictions, while recall (also called Sensitivity or True Positive Rate) is defined as
$Recall=\frac{TruePositives}{TruePositives+FalseNegatives}$
indicating the model’s capability to capture all unsafe cases (Yuan et al., 2024). By incorporating both these aspects, the F1 score serves as a robust indicator, especially in risk-sensitive applications where the accurate identification of unsafe instances is crucial.
Some related metrics are: specificity (also called True Negative Rate) is defined as:
$Specificity=\frac{TrueNegatives}{TrueNegatives+FalsePositives}$
and False Positive Rate (FPR) is calculated as:
$FPR=\frac{FalsePositives}{FalsePositives+TrueNegatives}$
also False Negative Rate (FNR) is measured as:
$FNR=\frac{FalseNegatives}{FalseNegatives+TruePositives}$
4
Refusal Rate (RR) measures the proportion of tasks an agent refuses to perform due to their aggressive, malicious, or otherwise unsafe nature (Zhang et al., 2024b; Andriushchenko et al., 2024). Defined as:
$Refusal~Rate=\frac{\#~Refused~Tasks}{\#~Total~Tasks}$
a higher RR indicates that the agent is more cautious and conservative, often rejecting potentially harmful or policy-violating requests. However, if the agent exhibits a high RR even on low-risk or benign tasks, it may be unnecessarily refusing requests. In contrast, a lower RR means the agent is more permissive, which can enhance user experience but might also lead to a greater risk of unsafe outcomes (Lee et al., 2024a). MLA-Trust (Yang et al., 2025c) instantiate RR as a Refusal-to-Execute Rate (RtE), where each output is judged (refuse / not refuse) by a specialized model such as GPT-4 or Longformer, following validated labeling protocols.
5 Leakage Rate (LR) evaluates how often an agent unintentionally leaks sensitive or private information. In PrivacyLens (Shao et al., 2024), a set $S$ of sensitive data is defined, and for each trajectory $\tau$, an agent output $a_{\tau}$ is considered a leakage event if any item in the sensitive data set $S$ can be inferred from it. Formally,
$LR=\frac{\#~Leakage~Cases}{\#~Total~Cases}$
AgentDAM (Zharmagambetov et al., 2025) likewise employs the LR metric to quantify any instances where sensitive data appears in an agent’s action outputs.
An adjusted version of the LR metric can further quantifies how often sensitive information might be leaked while also considering the helpfulness of the agent’s responses (Shao et al., 2024). It is defined as $LR_{h}=\frac{\#~Leakage~Cases~with~Positive~Helpfulness}{\#~Total~Cases~with~Positive~Helpfulness}$.
6 Cultural and Social Norms Metrics These metrics assess how well an agent recognizes and responds to cultural or social norms in user interactions:
Awareness Coverage Rate (AC-R) measures an agent’s ability to detect violations of cultural or social norms. A user query and the agent response are evaluated by an LLM judge, who determines whether the agent acknowledges any potential violation. The AC-R score is then computed as the proportion of queries where such violations are accurately recognized by agents (Qiu et al., 2025).
Educational Rate (Edu-R) measures whether the agent provides appropriate guidance or corrective feedback once a violation is detected. Automatic evaluator checks if the agent offers constructive advice in response to the identified norm breach. The Edu-R score is calculated as the fraction of queries where the agent successfully advises the user on how to align with cultural and social expectations (Qiu et al., 2025).
7 Inference Cost refers to the computational and monetary resources required to generate model outputs during deployment. In ShieldAgentBench (Chen et al., 2025e), it is quantified by the average number of closed-source LLM API queries and the end-to-end response time (in seconds) per sample, measured from the system runtime logs. Including inference cost alongside accuracy and robustness metrics offers a more complete view of an agent’s real-world performance and efficiency trade-offs.
8 Effectiveness assesses an agent’s ability to correctly identify and describe safety risks in interaction logs. Following the methodology of R-Judge (Yuan et al., 2024), an LLM-based evaluator assigns a graded relevance score to each risk scenario, comparing the agent’s risk analysis against a human-annotated reference. This metric directly reflects the agent’s ability to identify and address real safety concerns.
9 Attempt Rate (AR) captures how often an agent attempts to follow an adversarial instruction, even if it never finishes the harmful task. In both RedTeamCUA (Liao et al., 2025) and VPI-Bench (Cao et al., 2025) LLM judges scrutinize the trajectory to flag any onset of unsafe behavior—RedTeamCUA uses a single LLM to detect beginnings of harmful actions, while VPI-Bench employs a majority vote of three frontier LLMs to decide whether an attack was “attempted”. It is also termed Risk Goal Intention (RGI) with the same idea in (Yang et al., 2025b).
10 Toxicity Score (TS) estimates how likely a response contains toxic, offensive, or harmful content by assigning a scalar value. (Yang et al., 2025c)
5.4 Measurements
5.4.1 Rule-based Measurements
Rule-based measurement involves the use of predefined rules or algorithms to compute evaluation metrics without manual annotation or LLM intervention (Luo et al., 2025a). Typically implemented in code, these rules automatically assess agent behavior against fixed, deterministic criteria, making this approach suitable for well-defined and objective evaluation standards.
This method is widely adopted across existing agent safety benchmarks. For instance, ShieldAgent (Chen et al., 2025e) adopts this approach to directly compute evaluation metrics. TrustAgent (Hua et al., 2024) measures the overlap of action trajectories to assess goal alignment and safety compliance. AgentDojo (Debenedetti et al., 2024) and InjecAgent (Zhan et al., 2024) compute ASR variants using predefined criteria to capture attack effectiveness and resilience. PrivacyLens (Shao et al., 2024) uses binary (yes/no) rule-based judgments on privacy-sensitive prompts. Likewise, ST-WebAgentBench (Levy et al., 2024) applies programmatic functions to evaluate policy compliance via DOM and action traces. Both MobileSafetyBench (Lee et al., 2024a), Agent-SafetyBench (Zhang et al., 2024e) and WASP (Evtimov et al., 2025) rely on rule-based checks for task success, and even for harm prevention, while Agent Security Bench (ASB) (Zhang et al., 2024b) adopts rule-based ASR calculations to quantify attack impact. In addition, AgentHarm (Andriushchenko et al., 2024) also employs predefined rules to evaluate most simple tasks, thereby minimizing dependence on LLM-based grading. Furthermore, SafeArena (Tur et al., 2025) implements rule-based task-completion checks by matching agent outputs against predefined reference objects and uses the Agent Risk Assessment (ARIA) framework’s four hierarchical risk rules to quantify harmful-task outcomes. OpenAgentSafety (Vijayvargiya et al., 2025) implements Python-based evaluators that inspect the final environment state to detect unsafe outcomes. MLA-Trust (Yang et al., 2025c) adopts keywords matching method to automatically compute Refusal Rate. RAS-Eval (Fu et al., 2025) compares each agent’s tool-call sequence against a human-annotated reference sequence to automatically compute Task Completion, Incompletion, and Fail rates. RiOSWorld (Yang et al., 2025b) implement per-risk, rule-based evaluators that compare the final executable state to the expected unsafe outcome. Hijacking JARVIS (Liu et al., 2025a) evaluates the metrics with deterministic checks against task goals and labeled target elements with human annotated ground truth for the static set. AgentDAM (Zharmagambetov et al., 2025) also automatically inspects the final environment state against predefined success criteria to deterministically judge whether the agent completed its task. GUI-Robust (Yang et al., 2025a) evaluates agents by ensuring alignment with ground-truth trajectories throughout the task. Ultimately, VWA-Adv benchmark (Wu et al., 2024a) uses a rule-based approach to evaluate agent robustness, which models the agent as a directed graph and calculating adversarial influence along edges.
Rule-based approach can also be seen in some safety and security evaluation frameworks. For instance, the DoomArena framework (Boisvert et al., 2025) provides a modular library of scripted attack scenarios and built-in, deterministic checks on the final environment state, enabling automated detection of unsafe outcomes across a wide variety of web, OS, and multi-agent challenges.
While rule-based methods are efficient, automated, and reproducible—enabling scalable evaluation—they lack the flexibility to handle nuanced or context-dependent agent behaviors, making it hard to account for diverse unsafe attempts.
5.4.2 LLM-as-a-judge Measurements
LLM-based measurement leverages LLMs, such as general models like GPT-4 or fine-tuned models, to compute evaluation metrics based on natural language understanding, reasoning, and contextual judgment (Luo et al., 2025a). Unlike rule-based methods that rely on fixed logic, LLM-based approaches utilize the interpretive abilities of LLMs to handle complex and open-ended scenarios, making them ideal for tasks where deterministic rules fall short.
This approach is increasingly adopted in recent agent safety and capability benchmarks. For example, R-Judge (Yuan et al., 2024) uses an LLM-as-a-judge framework to score open-ended safety analyses, while TrustAgent (Hua et al., 2024) employs GPT-4 to assess both helpfulness and safety in agent outputs. ToolEmu (Ruan et al., 2023) uses automatic LLM evaluators to score safety and helpfulness for each trajectory. Both PrivacyLens (Shao et al., 2024) and AgentDAM (Zharmagambetov et al., 2025) applies a LLM-based evaluator to determine whether sensitive information is inferable from an agent’s action or output.
Similarly, BrowserART (Kumar et al., 2024) and AgentHarm (Andriushchenko et al., 2024) use GPT-4o to classify harmful behaviors and evaluate refusals. CASA (Qiu et al., 2025) adopts GPT-4o across metrics to assess cultural and social awareness, SafeArena (Tur et al., 2025) feeds GPT-4o each agent’s trajectory and metadata to assign one of the four ARIA risk levels, ASB (Zhang et al., 2024b) uses LLMs to evaluate whether agents properly refuse unsafe instructions, and MLA-Trust (Yang et al., 2025c) employs auto-classifiers to evaluate response toxicity and the misguided rate. Furthermore, OpenAgentSafety (Vijayvargiya et al., 2025) uses GPT-4.1 to label each trajectory into one of four predefined safety categories to capture unsafe intent that may not manifest in the final environment state. OS-Harm (Kuntz et al., 2025) also employs an LLM judge to decide task completion, label safety, and pinpoint the first unsafe step, with human annotations validating and confirming the LLM judge’s effectiveness. HAICOSYSTEM scores trajectories with LLM judges using scenario-specific checklists, assigning five risk dimensions on a 10-10 to 0 scale, plus Goal Completion and Tool-Use Efficiency on a 0 to $10$ scale.
Furthermore, RedTeamCUA (Liao et al., 2025), RiOSWorld (Yang et al., 2025b) and WASP (Evtimov et al., 2025) rely on LLM to identify trajectories where the agent shows intent to pursue the unsafe objective even without completing it. Rather than using a single LLM judge, VPI-Bench (Cao et al., 2025) employs a majority vote of three frontier models to decide whether a malicious instruction was attempted and whether it was completed. DoomArena framework (Boisvert et al., 2025) also supports plugging in LLM-based monitors to review intermediate agent thoughts and actions, flagging subtler breaches that its rule-based scripts might miss.
LLM-based methods are highly flexible and capable of capturing nuanced behaviors and contextual subtleties that rule-based systems often miss. However, they may suffer from variability across model versions, increased computational cost, and potential inconsistencies or bias in subjective judgments.
5.4.3 Manual Judge Measurements
Manual measurement involves human evaluators who directly assess the agent’s behavior or output. This method is indispensable in scenarios that require subjective judgment, nuanced contextual understanding, or complex reasoning that automated or model-based evaluators may struggle to capture accurately.
Despite its strengths in interpretability and accuracy for ambiguous cases, manual evaluation is labor-intensive, difficult to scale, and prone to individual bias. These limitations make it impractical for large-scale benchmarking and may limit its overall adoption in recent works.
Nevertheless, manual labels remain a valuable source of ground truth. For instance, R-Judge (Yuan et al., 2024) incorporates a human-labeled test set to assess the quality of LLM-generated safety analyses, using manual annotations as the gold standard to validate automated or LLM-based scoring methods. Similarly, ToolEmu (Ruan et al., 2023) employ human annotators to label emulation quality and agent safety/helpfulness, providing a reference set to validate the LLM judges. In addition, SafeArena (Tur et al., 2025) conduct manual assessments of each trajectory using the ARIA framework.
6 Discussion
In the preceding sections, we have examined the current landscape of attacks, defenses, and evaluation methodologies pertinent to the security of CUAs. Building upon these insights, this discussion synthesizes key findings and outlines promising avenues for future research.
6.1 Key Insights
The rapid adoption of CUAs across diverse domains has revealed several pivotal observations:
  • Real-Time and Multimodal Emphasis: Unlike traditional LLM-based agents that primarily handle static text input, CUAs often operate in dynamic environments and interact with multiple input modalities, such as touch-based GUIs, images, and voice commands. This dual emphasis on real-time responsiveness and multimodal task comprehension introduces unique challenges such as handling long reasoning gaps, preventing multimodal hallucinations, and managing on-device resource constraints (Zhang and Zhang, 2023; Nong et al., 2024; Zhang et al., 2023).
  • UI Understanding Difficulties: Current benchmarking efforts reveal that many CUAs demonstrate suboptimal safety performance, reflecting gaps in robustness and risk awareness (Zhang et al., 2024e; Andriushchenko et al., 2024; Lee et al., 2024a; Zhang et al., 2024b). A significant portion of this shortfall arises from immature grounding techniques, which hinder agents’ ability to reliably interpret multimodal perceptions, especially vision tasks (Zheng et al., 2024; Zhang and Zhang, 2023; Liu et al., 2025b). These shortcomings underscore the need for more holistic training and test scenarios that address diverse threat models.
  • Limited Experimental Scenarios: Many CUAs are tested in highly constrained settings that fail to capture the breadth of real-world tasks. For instance, the action space is restricted by excluding multi-touch or irregular gestures in (Zhang et al., 2023), while (Zhang and Zhang, 2023) focuses on small-scale GUI agents in a single environment (AITW), limiting broader investigations. Additionally, PC-Agent has only explored productivity scenarios, leaving potential social and entertainment applications largely unexamined (Liu et al., 2025b).
  • Transparency Deficits: A number of CUA providers neither publish safety policies nor disclose systematic evaluation outcomes, making it difficult for users and policymakers to assess an agent’s reliability. The absence of transparent risk disclosures impedes accountability mechanisms and could enable unchecked vulnerabilities (Shi et al., 2024; Hua et al., 2024; Hu et al., 2024).
6.2 Future Directions
Tackling these challenges requires a multifaceted research agenda, integrating both technical innovations and governance considerations:
  • Integrated Defense Mechanisms: Research on robust defenses spans active adversarial mitigation, environmental verification, and backdoor detection. Proposed methods include integrating modules for trustworthiness checks and leveraging multi-agent approaches for role-specific security tasks (Chen et al., 2025d; Zeng et al., 2024; Tian et al., 2023).
  • Real-time Comprehensive Benchmarking: Broader and more dynamic benchmarks are essential for capturing real-world complexity. Future evaluations should incorporate tasks requiring advanced domain expertise, testing agents’ resilience under challenging conditions and adaptive attacks (Zhang et al., 2024e; Levy et al., 2024; Debenedetti et al., 2024; Andriushchenko et al., 2024).
  • Transparency and Accountability: Establishing standardized guidelines for disclosing safety policies and reporting evaluation protocols can strengthen trust in CUAs. Such measures could include enforced policy publication, structured reporting of risk assessments, and independent audits (Hua et al., 2024; Shi et al., 2024; Shao et al., 2024).
  • Human-Agent Collaboration: Incorporating mechanisms for human oversight—especially in high-risk domains—can mitigate the potential harm of fully autonomous operations. Systems designed to allow timely human intervention and clear explanations of agent decisions will improve safety and foster user confidence (Wang et al., 2023; Fang et al., 2024a; Sager et al., 2025).
By advancing defense strategies, refining benchmarks, promoting transparency, and integrating principled human oversight, researchers and developers can elevate both the reliability and trustworthiness of CUAs. Addressing these multifaceted challenges will be central to ensuring that these agents are not only effective in diverse applications but also safe to deploy in real-world environments.
7 Conclusion
The rapid advancement of Computer-Using Agents (CUAs) has introduced powerful capabilities for GUI automation, but also significant safety challenges. In this survey, we have presented a comprehensive examination of these challenges, systematically analyzing risks across four key dimensions: defining CUAs and their components, categorizing both intrinsic and extrinsic threats, evaluating defense strategies, and reviewing benchmarking approaches.
Looking ahead, three priorities emerge: (1) the development of unified safety standards applicable to various CUA implementations, (2) the creation of robust testing environments that accurately simulate real-world complexities, and (3) the enhancement of transparency to foster user trust. Achieving these objectives will require interdisciplinary collaboration, integrating insights from AI, security, and human-computer interaction.
As CUAs become increasingly embedded in critical systems, their safety can no longer be an afterthought. This survey provides a foundation for future research, emphasizing that security and capability must advance together. Future work should focus on creating more resilient agents while establishing frameworks for responsible deployment, ensuring these powerful tools benefit users without introducing new risks.
Appendix A Complete Taxonomy
{forest}
for tree= forked edges, draw, rounded corners, node options=align=center, s sep=12pt, calign=center, grow=east, reversed=true, parent anchor=east, child anchor=north, font=, , where level=1text width=100pt, fill=blue!10, rotate=90, where level=2text width=100pt, fill=pink!30, child anchor=west, where level=3text width=50pt, fill=yellow!30, child anchor=west, where level=4text width=100pt, fill=cyan!20, child anchor=west, where level=5child anchor=west, [Computer-Using Agent Safety, fill=gray!20, rotate=90 [Threats [Intrinsic Threats [Perception [UI Understanding and Grounding Difficulties, text width=150pt [Chen et al. (2025c), Pahuja et al. (2025), Nong et al. (2024), Zheng et al. (2024), Liu et al. (2025b) , text width = 350pt] ] ] [Brain [Scheduling Errors, text width=150pt [Zhang and Zhang (2023), Deng et al. (2024) ] ] [Misalignment, text width=150pt [Ma et al. (2024) ] ] [Hallucination, text width=150pt [Deng et al. (2024) ] ] [Excessive Context Length, text width=150pt [Zhang and Zhang (2023), Yang et al. (2024a) ] ] [Social and Cultural Concerns, text width=150pt [Qiu et al. (2025) ] ] [Response Latency, text width=150pt [Wen et al. (2023), Zhang and Zhang (2023), Li et al. (2020), Li et al. (2021) ] ] ] [Action [API Call Error, text width=150pt [Deng et al. (2024), Nong et al. (2024), Zhang and Zhang (2023) ] ] ] ] [Extrinsic Threats [Adversarial Attack, text width=150pt, fill=cyan!20 [Wu et al. (2024a), Ma et al. (2024), Zhang et al. (2024c), Aichberger et al. (2025), Zhao et al. (2025), Wu et al. (2025) , text width=300pt, fill=white] ] [Prompt Injection Attack, text width=150pt, fill=cyan!20 [Zhan et al. (2024), Liao et al. (2024), Wu et al. (2024b), Xu et al. (2024), Debenedetti et al. (2024), Mudryi et al. (2025), Liu et al. (2023b), Lupinacci et al. (2025), Kuntz et al. (2025), Liao et al. (2025), Evtimov et al. (2025), Wu et al. (2025), Cao et al. (2025), Shapira et al. (2025), Chen et al. (2025a), Liu et al. (2025a), Wang et al. (2025a), Nakash et al. (2024), Johnson et al. (2025), Wang et al. (2025e), Zhan et al. (2025), Lu et al. (2025), Wang et al. (2025g), Wang et al. (2025c) , text width=400pt, fill=white] ] [Jailbreak, text width =150pt, fill=cyan!20 [Mo et al. (2024), Chu et al. (2024), Zou et al. (2023), Liu et al. (2023a), Kumar et al. (2024), Tian et al. (2023), Zhang et al. (2024d), Mao et al. (2025), Kuntz et al. (2025), Gu et al. (2024), Qi et al. (2025) , text width=450pt, fill=white] ] [Memory injection Attack, text width =150pt, fill=cyan!20 [Patlan et al. (2025a), Patlan et al. (2025b), text width=400pt, fill=white] ] [Backdoor Attack, text width =150pt, fill=cyan!20 [Yang et al. (2024b), Wang et al. (2024), Zhu et al. (2025b), Chen et al. (2024b), Boisvert et al. , Ye et al. (2025), Wang et al. (2025f), Cheng et al. (2025), Lupinacci et al. (2025) , text width = 400pt, fill=white] ] [Reasoning Gap Attack, text width = 150pt, fill=cyan!20 [Chen et al. (2025d) , fill=white] ] [System Sabotage, text width =150pt, fill=cyan!20 [Luo et al. (2025b) , fill=white] ] [Web Hacking, text width =150pt, fill=cyan!20 [(Fang et al., 2024b), fill=white] ] ] ] [Defenses [Environmental Constraints, text width=210pt, fill=cyan!20 [Yang et al. (2024c), Nong et al. (2024), Zhang et al. (2023), Mahadevan et al. (2025), Huang et al. (2025), text width=400pt, fill=white] ] [Input Validation, text width=210pt, fill=cyan!20 [Kumar et al. (2024), Wen et al. (2023), Zhang et al. (2024c), Tshimula et al. (2024), Shi et al. (2025a), Zhong et al. (2025), Ferrag et al. (2025), text width=450pt, fill=white] ] [Defensive Prompting, text width=210pt, fill=cyan!20 [Debenedetti et al. (2024), Zhang et al. (2024c), Wu et al. (2024a), Chen et al. (2025d), text width=450pt, fill=white] ] [Data Sanitization,text width=210pt, fill=cyan!20 [Yang et al. (2024b), Jones et al. (2025), Wang et al. (2025b),text width=300pt, fill=white] ] [Adversarial Training,text width=210pt, fill=cyan!20 [ Yu et al. (2020), Wu et al. (2024a), Zhou et al. (2025), Yu et al. (2025),text width=400pt, fill=white] ] [Output Monitoring, text width=210pt, fill=cyan!20 [Fang et al. (2024a),Wen et al. (2023), Hua et al. (2024), Lee et al. (2025), Shi et al. (2025b), text width=500pt, fill=white] ] [Model Inspection, text width=210pt, fill=cyan!20 [Wang et al. (2025d), Yang et al. (2024b), Huang et al. (2025), Zhu et al. (2025b), text width=400pt, fill=white] ] [Cross Verification, text width=210pt, fill=cyan!20 [Zeng et al. (2024), Huang et al. (2024), Yang et al. (2024a), Xiang et al. (2024), Luo et al. (2025b), Barua et al. (2025), Zhu et al. (2025a), Li et al. (2025), Fan and Li (2025), text width=500pt, fill=white] ] [Continuous Learning and Adaptation, text width=210pt, fill=cyan!20 [Tian et al. (2023), Ma et al. (2024), Luo et al. (2025b), Zhan et al. (2025), Zhang et al. (2025b), text width=500pt, fill=white] ] [Transparentize, text width=210pt, fill=cyan!20 [Hu et al. (2024), Sager et al. (2025), Chen et al. (2025b), text width=300pt, fill=white] ] [Topology-Guided, text width=210pt, fill=cyan!20 [Wang et al. (2025d), text width=100pt, fill=white] ] [Perception Algorithms Synergy, text width=210pt, fill=cyan!20 [Zheng et al. (2024), Nong et al. (2024), Liu et al. (2025b), Yang et al. (2024a), text width=400pt, fill=white] ] [Planning-Centric Architecture Refinement, text width=210pt, fill=cyan!20 [Zhang and Zhang (2023), Deng et al. (2024), Wen et al. (2023), Liu et al. (2025b), text width=420pt, fill=white] ] [Pre-defined Regulatory Compliance, text width=210pt, fill=cyan!20 [Zhang et al. (2024e), Mahadevan et al. (2025), Chen et al. (2025e), Zhang et al. (2025a), text width=450pt, fill=white] ] ] ]
Figure 1: A comprehensive taxonomy of Computer-Using Agent threats and defences.